### v1

# Nonlinear Causality in Brain Networks: With Application to Motor Imagery vs Execution

Abstract
One fundamental challenge of data-driven analysis in neuroscience is modeling causal interactions and exploring the connectivity of nodes in a brain network. To this end, many statistical methods, relying on various perspectives and employing different data modalities, are being developed to examine and comprehend the underlying causal structures inherent to brain dynamics. This study develops a novel statistical approach enabling neuroscientists to dissect causal interactions in multichannel EEG recordings. Our approach, TAR4C, uses the threshold autoregressive model to describe the causal interaction between nodes or clusters of nodes in a brain network. The perspective involves testing whether one node, which may represent a brain region, can control the dynamics of the other. The node that has such an impact on the other is called a threshold variable and can be classified as a causative because its functionality is the leading source operating as an instantaneous switching mechanism that regulates the time-varying autoregressive structure of the other. This statistical concept is commonly referred to as threshold non-linearity. Once threshold non-linearity has been verified between a pair of nodes, the subsequent essential facet of TAR modeling is to assess the predictive ability of the causal node for the current activity on the other and represent causal interactions in autoregressive terms. This predictive ability is what underlies Granger causality. The TAR4C approach can discover non-linear and time-dependent causal interactions without negating the G-causality perspective. The efficacy of the proposed approach is exemplified by analyzing the EEG signals recorded during the motor movement/imagery experiment. The similarities and differences between the causal interactions manifesting during the execution and the imagery of a given motor movement are demonstrated by analyzing EEG recordings from multiple subjects.
Keywords: Regime Switching Time Series Models; Threshold AR; Nonlinear Causality; Time-varying Causality; Brain Connectivity
1 Introduction
The exploration and understanding of the complex connectivity and causality intrinsic to networks of time series data represents a pivotal challenge that necessitates significant methodological enhancements, particularly within the domains of neuroscience and statistical analyses for brain imaging data (Hamedi et al., 2016; Sheikhattar et al., 2018; Farahani et al., 2019; Runge et al., 2019; Biswas and Shlizerman, 2022; Varley and Sporns, 2022). As our understanding of brain function deepens, it becomes increasingly evident that brain connectivity is both dynamic and nonlinear, demanding sophisticated analytical approaches to decode the underlying interactions (Razi and Friston, 2016; Cao et al., 2022; Siddiqi et al., 2022). The further progress and analytical capabilities attained for this objective are also propelling progress in other fields, including economics and climate science, when and where the objective is to analyze time series data from a network/graph perspective (Papana et al., 2017; Gong et al., 2019; Upadhyay et al., 2020; Wang and Huang, 2021; Nowack et al., 2020; Naghipour et al., 2021; Holme and Rocha, 2023; Tjøstheim et al., 2023).
This study aims to propose a conceptually comprehensive methodology that captures nonlinear causal connectivities between any given Nodes of Interest (NOIs) within a network of time series. To align with this aim, a data analytical objective is set to explore causal interactions within a network of electroencephalogram (EEG) time series data. While conventional and state-of-the-art methods are undoubtedly practical and helpful, they are limited in their ability to incorporate the nonlinear, especially dynamic phase-shifting (i.e., phase transition) and time-varying nature of these interactions (Jirsa et al., 1994; Beggs and Plenz, 2003; Beggs, 2008, 2022; Corsi et al., 2024; Shriki et al., 2013; Sporns, 2016; Bocaccio and Tagliazucchi, 2024).
1.1 Literature Background Specific to Causal Inference in EEG
A significant number of methodologies, including Granger Causality (GC) (Ding et al., 2006; Marinazzo et al., 2011; Barrett et al., 2012; Zhao et al., 2013; Seth et al., 2015; Benozzo et al., 2017; Stokes and Purdon, 2017; Sheikhattar et al., 2018; Dhamala et al., 2018; Zhang et al., 2019; Wismüller et al., 2021; Heyse et al., 2021; Biswas and Ombao, 2022), Transfer Entropy (TE) (Vicente et al., 2011; Huang et al., 2015; Diks and Fang, 2017; Ursino et al., 2020; Redondo et al., 2023), Directed Transfer Function (DTF) (Babiloni et al., 2005; Wilke et al., 2008; Wang et al., 2016; Heyse et al., 2021; Li and Zhang, 2022), Partial Directed Coherence (PDC) (Chiang et al., 2009; Varotto et al., 2012; Omidvarnia et al., 2013; He et al., 2014; Mill et al., 2017; Gaxiola-Tirado et al., 2017), Dynamic Causal Modeling (DCM) (Friston et al., 2003; Kiebel et al., 2008; Chen et al., 2008; Daunizeau et al., 2009; Chicharro and Panzeri, 2014; Van de Steen et al., 2019; Wang and Liang, 2021), Bayesian modeling (Prado et al., 2001; Rajapakse and Zhou, 2007; Gorrostieta et al., 2012, 2013; Yu et al., 2016; Chiang et al., 2017; Kook et al., 2020; Hu et al., 2020; Granados-Garcia et al., 2024), and Machine / Deep Learning (ML / DL) modeling (Schirrmeister et al., 2017; Chai et al., 2019; Phang et al., 2019; Zhang et al., 2019, 2021; Gao et al., 2021; Rosol et al., 2022; Luo et al., 2022; Liu et al., 2023; Bocaccio and Tagliazucchi, 2024) based approaches, have been proposed as means of inferring causality and connectivity in EEG, with varying degrees of focus on specific objective aspects.
Furthermore, approach-based literature surveys outlined by Haufe et al. (2011); Razi and Friston (2016); Hamedi et al. (2016); Farahani et al. (2019); Ombao and Pinto (2022); Siddiqi et al. (2022); Shojaie and Fox (2022); Biswas and Shlizerman (2022); Chiarion et al. (2023); Zheng et al. (2023) offer valuable insights into the current framework on human brain connectivity.
The conventional GC method employs linear autoregressive models, which assess whether the past values of one time series can enhance the prediction of another. It should be noted that due to the inherent linearity and Gaussianity assumption of conventional GC, its usage in the complex, nonlinear dynamics of brain signals will lead to limited applications and interpretations (Lusch et al., 2016; Shojaie and Fox, 2022). In light of these structural restrictions, researchers have sought to develop alternative approaches that would permit the G causality framework to be extended to the nonlinear dynamics characterizing real-world contexts (e.g., Marinazzo et al., 2011; Zhao et al., 2013; Henderson and Michailidis, 2014; Chen et al., 2017; Runge et al., 2019; Wismüller et al., 2021; Tank et al., 2022). Developing methodology for Nonlinear GC inference remains an active area of research due to several inherent challenges. These include the difficulty of determining the degree of nonlinearity, the prevalence of overfitting issues, the incompatibility of the method with high dimensionality, the generation of high variance estimators, the necessity for a large number of hyperparameters, the use of black-box routines, and the complexity of interpreting the results.
On the other hand, TE is a nonparametric, model-free method for quantifying directional dependencies between time series, particularly useful for capturing nonlinear interactions. However, TE faces challenges related to computational complexity, data requirements, and interpretability. It requires large datasets for reliable estimates, and its sensitivity to noise and data sparsity, especially in EEG, can lead to false positives or spurious connections. TE’s performance is further influenced by parameter choices, such as embedding delay and dimension, adding complexity. In multivariate systems, TE may mistakenly attribute indirect connections as direct, leading to inaccurate conclusions. Partial Transfer Entropy (PTE) addresses this issue by considering other variables but retains many of TE’s computational and interpretative challenges. Despite these limitations, TE’s ability to reveal hidden causal relationships makes it a powerful tool across disciplines, though ongoing refinement is needed to improve its practicality and clarity (Wibral et al., 2014; Ekhlasi et al., 2023; Das and Porfiri, 2023).
Dynamic Causal Modeling (DCM) is a Bayesian framework used in neuroimaging to model how brain regions influence each other based on neural and hemodynamic principles. Unlike Granger causality, DCM uses mechanistic differential equations to describe neural activity over time and integrates prior knowledge about brain structure to estimate causal influences. While these biologically grounded priors allow for detailed inferences about brain dynamics, they also risk model misspecification if the assumptions are incorrect, leading to biased results. Additionally, DCM’s computational intensity and sensitivity to the selection of priors make model comparison challenging, as different priors can lead to varying conclusions. Despite these limitations, DCM is valued for its ability to incorporate biophysical constraints, making it particularly useful for testing specific hypotheses about neural circuits, such as the directionality of connections during cognitive tasks (Weichwald and Peters, 2021; Lohmann et al., 2012; Razi and Friston, 2016).
In contrast, machine learning models have gained popularity for their ability to detect complex, nonlinear relationships in EEG data, but they are often criticized for their lack of interpretability. These models can effectively classify and predict brain states by recognizing subtle patterns in high-dimensional data. However, they often operate as ”black boxes,” offering little insight into the underlying causal mechanisms of neural activity. This ambiguity can impede the extraction of meaningful insights about causal relationships in brain networks. Moreover, machine-learning approaches require large, well-labeled datasets, which are often scarce in neuroscience. EEG data are particularly noisy, and machine learning models are highly sensitive to this noise, which can lead to poor generalization across different subjects, experimental conditions, or datasets if the overfitting issue is not properly addressed. Furthermore, while ML methods are adept at capturing associations and providing strong predictive power, distinguishing between causality and mere correlation remains elusive, requiring the integration of additional causal reasoning frameworks. Reproducibility across studies can also be a concern, as model assumptions, dataset variability, and preprocessing differences can result in inconsistent findings across studies (Wein et al., 2021; Ji et al., 2021). Despite these challenges, ML’s flexibility and capacity for handling vast amounts of data offer exciting prospects, especially when combined with rigorous causal inference methods.
1.2 Challenges of Inferring Causal Interactions in EEG Data
Inferring causal interactions in brain activity is challenging due to the brain’s complex, nonlinear, and time-varying dynamics. Connectivity constantly changes with stimuli, cognitive processes, and internal conditions, making it difficult to separate direct causal relationships from correlations and interdependence.
It is well-documented that brain functioning is characterized by various types of nonlinearities (Mayer-Kress, 1998; Breakspear and Terry, 2002; Breakspear, 2017; Lynn and Bassett, 2019; O’Donnell, 2023). Neural nonlinearities are complex, non-proportional relationships in brain function that appear in various forms (e.g., synaptic plasticity, neural oscillations, neuronal firing rates). The literature on the exploration of brain connectivity has sought to encompass the phenomenon of nonlinearity through the use of state-of-the-art approaches and empirical investigations. However, the notion that the brain functions in a non-linear way is taken as an umbrella concept, and it is assumed that each study addresses nonlinearity appropriately. Since nonlinearity can take various forms, it may be more accurate to refer to ’nonlinearities’ rather than ’nonlinearity’ in brain function. A single methodology cannot capture all these variations or support unified conclusions. This paper analyzes the abrupt phase-shifting dynamics of the brain from an unexplored nonlinearity perspective in relation to EEG data, aligning with findings that the brain processes information with nonlinear, abrupt phase shifts (Beggs, 2008; Razi and Friston, 2016).
A further challenge in inferring causality from EEG data resides in the limited capacity of EEG to depict spatial resolution of neuronal activities in sufficient detail. This insufficiency makes it difficult to localize neural sources, complicating the task of identifying genuine causal relationships between brain regions. The recorded signals often reflect aggregated activity from multiple sources, distorted by volume conduction, which can create undecomposable correlations and spurious causal inferences, especially for nearby channels. While methods like source localization and beamforming can help mitigate these issues, they also introduce their own limitations and assumptions that must be carefully considered. The causal inference obtained from EEG often reflects a simplified model of brain connectivity, lacking the granularity needed to understand neural mechanisms comprehensively. While EEG reveals the directional flow of information between channels, it offers limited insight into the biophysical processes involved, such as synaptic or neurotransmitter activity. Therefore, EEG-based causality should be interpreted cautiously, as it represents statistical connections rather than direct neural interactions. Despite its significant temporal resolution, EEG’s limited spatial and mechanistic detail restricts its value in brain connectivity studies. However, it remains useful for exploring neural timing and sequence, especially when combined with other methods (Bastos and Schoffelen, 2016).
1.3 Objective of the Study
This study aims to introduce a new approach, the Threshold Autoregressive model for Causality (TAR4C), to discover nonlinear and time-varying causal interactions, especially when the goal is to dissect the connectivity between only a few nodes in multi-node networks of time series, focusing on EEG data. The specific objectives are as follows:
Introducing Threshold Nonlinearity for Connectivity Inference: Existing methods overlook a key brain principle: abrupt transition dynamics, often focusing on specific aspects of nonlinearity. In the TAR framework, the threshold variable (i.e., node/channel) triggers activity in another node when its nonlinearity effect is significant. The threshold variable’s activity governs the other, suggesting a connection between the two nodes, even without additional evidence. This approach reveals phase transitions through threshold variables, offering a new understanding of EEG connectivity and dynamics.
Modeling Nonlinear and Time-Varying Causal Interactions: Another goal is to model how different nodes in a network influence each other under different conditions by validating nonlinear and time-varying interactions. Applying TAR4C to EEG data enables us to study the causal relationships as threshold variable’s states change, particularly during motor tasks like execution and imagery. TAR4C is, therefore, capable of tracking causal connectivity alterations over time, which is not addressed by current methodologies.
Comparing Connectivity Patterns in Motor Execution and Imagery: The study also aims to examine differences in brain connectivity during motor execution (MM) and motor imagery (MI). It investigates how causal interactions between regions, such as the prefrontal and motor cortices, vary across tasks involving different gestures (e.g., left fist, right fist, both fists, both feet). By applying TAR4C, the study discovers distinct patterns in how the brain responds to real and imagined motor tasks.
Quantifying Connectivity with New Indices: The TAR4C method introduces two new metrics: the Threshold Connectivity Index (TCI) and the Threshold Granger Causality Index (TGCI). These indices measure the strength and direction of causal interactions between EEG channels (e.g., nodes). The TCI identifies threshold-based connectivity, while the TGCI evaluates the predictive relationships between channels. These indices provide a simplified representation of nonlinear brain interactions.
Expanding TAR4C Beyond EEG Networks: Although focused on EEG data, the TAR4C framework can be applied to other time series networks, like those in economics or climate, when the goal is to uncover relationships between specific variables in high-dimensional settings. This study positions TAR4C as a general tool for detecting nonlinear causal interactions across various fields, emphasizing the broader relevance of threshold autoregressive models in understanding complex connectivity patterns.
2 Threshold Autoregressive Modeling for Exploring Connectivity and Causality
This study proposes an approach based on threshold autoregressive (TAR) modeling to explore connectivity within time series networks and a novel application to multi-channel motor EEG recordings presented as a case study. The brain, a highly dynamic structure, constantly shifts between different connectivity states in response to stimuli and tasks. EEG devices measure brain activity through signals captured by electrodes (channels/nodes) on the scalp. These signals, rich in temporal information, are valuable data for studying brain connectivity during cognitive processes, neurological disorders, and responses to various stimuli.
EEG recordings reflect time-varying connectivities and phase transitions that occur during brain activities, such as perception, cognition, and motor tasks. This study assumes that these transitions are governed by phase-shifting mechanisms that form effective connections between brain regions. The TAR model, designed to handle regime shifts, can track these transitions by specifying functionals that control the dynamics. By identifying these phase changes, the TAR model provides a practical tool for discovering causal interactions among EEG channels, offering insights into how different channels influence one another.
2.1 Inference in high-dimensional networks of time series
Most phenomena in the real world are composed of interconnected components. Temporal observations of multi-component systems form multivariate time series datasets. In the context of this study, it is appropriate to refer to this data structure as a network of time series, as this terminology is more inclusive and clearly emphasizes interconnectedness. Statistical inference in such high-dimensional networks of time series continues to pose a significant analytical challenge. For instance, in an Economics study designed to assess the economic conditions of a country or countries, it will be essential to analyze datasets of time-dependent indicators. In Finance, the problem of identifying the best asset allocation among all possible stocks would require analyzing a complex network of time series. Or, for example, to study climate change, it would be necessary to work on a network of time series of different meteorological variables from many spatial locations. Ombao and Pinto (2022) equip a unified framework with statistical methods for addressing such networks of time series based problems.
In such interconnected multi-node networks, if the primary purpose and scope of the research are to determine the connectivity between any specific two nodes, the confounding effects of the other nodes in the network cannot be ignored. In this case, a multi-parameter model to analyze the whole network or a method to eliminate the effects of other nodes in the network should be followed. Since the estimation and interpretation of multi-parameter models become more difficult as the number of parameters increases, instead of using a multi-parameter model, we adopt a practical and innovative dimensionality reduction approach proposed by Aslan and Ombao (2024) to determine the connectivity between two specific nodes. This approach employs spectral dynamic principal component analysis (sDPCA) in the frequency domain to create a low-dimensional representation of all other nodes. The resulting principal component scores are subsequently removed from the nodes of interest, thus eliminating the confounding effects. This method maintains the temporal and spectral characteristics of the data, which is crucial for accurate connectivity analysis. It ensures that the effects of other nodes are effectively isolated, allowing for a precise determination of connectivity between the two nodes.
Define $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ to be a network of time series with $n$ nodes. In this paper, we consider a network representing brain activity through EEG recordings, where each component (i.e., node or channel) is a time series. Within this network, denoted as $N$, $Y^{*}$, and $X^{*}$ represent any two nodes of interest. For problems that require component-wise solutions, our objective is to investigate the interaction or effective connectivity between any pair of nodes rather than the relationships within the entire network. These two nodes are influenced by other nodes due to the network’s interconnected structure. Therefore, to accurately analyze the effective connection between any two nodes of interest, it is necessary to isolate them from the influences of other nodes. As demonstrated and proposed by Aslan and Ombao (2024), sDPCA in the frequency domain offers practical benefits and is highly effective for the alleviation of dimensionality problems in high-dimensional settings. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
$Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
$X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
where $f_{z}(.)$ is a function that accounts for the dynamic variability of the excluded nodes. In practice, $f_{z}(.)$ serves as a filter and can take any form, provided it effectively summarizes the net discharge of the excluded nodes. See Brillinger (1974); Ombao and Ho (2006); Hörmann et al. (2015); Hörmann and Kidzinski (2024); Wang et al. (2019); Ombao and Pinto (2022); Aslan and Ombao (2024) for more details.
2.2 Background on the threshold autoregressive model
The concept of switching between regimes based on a change point was first introduced by Quandt (1958) in linear regression. Two decades later, Tong (1978) and Tong and Lim (1980) extended this idea to nonlinear time-series analysis through threshold non-linearity. Since then, the threshold autoregressive (TAR) model has been applied widely across various domains. As detailed by Tong (1990), TAR models offer practical methodologies for exploring complex dynamical systems in time series, as they capture nonlinear dynamics, limit cycles, abrupt jumps, and asymmetries.
The primary goal of the TAR model is to identify threshold-driven regimes in time-varying systems and provide autoregressive representations of these regimes. The thresholding mechanism makes the TAR model particularly useful, as it decomposes complex stochastic systems into simpler subsystems, allowing for more efficient implementation and better interpretation. TAR models are well-suited to represent systems with shifting levels and sudden bursts, often seen in time series data such as EEG signals. For a comprehensive review of the applications of threshold nonlinearity and TAR models, we refer the reader to Hansen (2011) and Chen et al. (2011). In this study, we introduce the TAR model as an effective framework for analyzing the complex and highly dynamic brain activity captured through EEG signals.
For a stochastic process denoted as, $Y_{t}$, a general TAR model having $k$ regimes can be represented as follows:
$Y_{t}=\sum_{j=1}^{k}\left[(\phi_{0}^{(j)}+\sum_{i=1}^{p_{j}}\phi_{i}^{(j)}Y_{t-i})\ I(r_{j-1}<Z_{t-d}\leq r_{j})\right]+\ \varepsilon_{t},$ (1)
[…]
For example, if at time $t$, let $Z_{t-d}=a\in(r_{j-1},r_{j})$ then the active regime at that time is characterized by linear AR model;
$Y_{t}=\phi_{0}^{(j)}+\sum_{i=1}^{p_{j}}\phi_{i}^{(j)}Y_{t-i}\ +\ \varepsilon_{t}^{(j)}.$
The overall process $Y_{t}$ is nonlinear if there are at least two regimes with different models, even if each is linear. Therefore, by describing the different regimes by means of threshold variables, we can ascertain the temporal behavior of the dynamics.
The unknown parameters for the model in Equation (1) can be assembled as follows:
$\Omega=\left[(\phi_{0}^{(1)},\phi_{i:p_{1}}^{(1)}),\ldots,(\phi_{0}^{(k)},\phi_{i:p_{k}}^{(k)})\right].$
The hyperparameters of the model can be shown as;
$\Gamma=\{k,p_{j},d,r\in[\text{min}(Z_{t-d}),\text{max}(Z_{t-d})]\}.$
Since the TAR model is a piece-wise linear model, the estimates of the parameters can be readily obtained by using the least squares (LS) method conditional on hyperparameters (i.e., estimation is equivalent to the maximum likelihood estimation under the auxiliary assumption that $\varepsilon_{t}$ is i.i.d.$N(0,\sigma_{\varepsilon_{t}}^{2})$). Thus, the conditional least squares (CLS) method has a significant advantage in that it allows the exploration of the threshold parameter and hyperparameters without imposing any restrictions. Chan (1993) demonstrated the consistency of the CLS method under certain mild conditions, such as stationarity and the presence of a discontinuous conditional mean function.
The LS estimators $\widehat{\Omega}$, $\widehat{\Gamma}$ jointly minimize the following objective function:
$\mathcal{S}({\Omega},{\Gamma})=\sum_{t=1}^{T}\left[Y_{t}-\sum_{j=1}^{k}\left[(\phi_{0}^{(j)}+\sum_{i=1}^{p_{j}}\phi_{i}^{(j)}Y_{t-i})I(r_{j-1}<Z_{t-d}\leq r_{j})\right]\right]^{2},$ (2)
$\widehat{\Omega},\widehat{\Gamma}=\operatorname*{\arg\min}_{\Omega,\Gamma}\mathcal{S}({\Omega},{\Gamma})$ (3)
where $T$ denotes the sample size. The LS estimators $\widehat{\Omega}$ and the minimization of the objective function in Equation (2) can be achieved conditionally by a grid search considering all viable combinations of hyperparameters $\{j,p_{j},r,d|j=1,\ldots,k\}$. The optimization requires the execution of approximately $T^{k-1}/2\times d\times p_{(1)}\times\ldots\times p_{(k)}$ arranged autoregressions. Arranged autoregression is an autoregression where cases are rearranged by $\Gamma$-dependent sample splitting. Note that rearrangement of the sample based on the threshold values does not change the temporal structure of the series within cases (Tsay, 1989).
The asymptotic theory of LS estimation and inference for TAR models was addressed and developed by Chan (1993), Chan and Tsay (1998), Hansen (1997, 2000), Li and Ling (2012), Zhang et al. (2023).
Remark. The model in Equation (1) is called self-exciting threshold autoregressive (SETAR) model if the threshold variable is the same as the actual time series, i.e., $Z_{t-d}=Y_{t-d}$. In this case, the regime of the time series $Y_{t}$ would be determined by its own past value, $Y_{t-d}$.
2.3 Proposed Approach: TAR for causality in EEG networks
EEG functions as a network of interconnected components, where connectivity between node pairs can be the subject of investigation. This study proposes a 2-regime TAR model based approach to dissect causal interactions between EEG channels. The main objective is to map the net causal influence each node receives from others through pairwise connectivity analysis while minimizing interference from excluded nodes.
Let $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ represents $n$-dimensional network of multiple EEG time series; $Y^{*}$ and $X^{*}$ be the nodes of interest in this network, between which we aim to discover the causal relationship. Considering $Y$ and $X$ be the neutralized counterparts of $Y^{*}$ and $X^{*}$ from the impact of other nodes in the network, the ”net” activities of nodes $Y^{*}$ and $X^{*}$ at time $t$ can be denoted as $Y_{t}$ and $X_{t}$. Thereby, the 2-regime TAR model for the node $Y$, and the node $X$, which is treated as both the threshold and predictive variable, can be formulated as follows:
$Y_{t}=\begin{cases}\phi_{0}^{(1)}+\sum_{i=1}^{p_{1}}\phi_{i}^{(1)}Y_{t-i}+\sum_{\ell=1}^{q_{1}}\theta_{\ell}^{(1)}X_{t-\ell}\ +\ \varepsilon_{t},&\text{if }u(X_{t-d})\leq r\\ \phi_{0}^{(2)}+\sum_{i=1}^{p_{2}}\phi_{i}^{(1)}Y_{t-i}+\sum_{\ell=1}^{q_{2}}\theta_{\ell}^{(2)}X_{t-\ell}\ +\ \varepsilon_{t},&\text{if }u(X_{t-d})>r\end{cases},$ (4)
where $u(X_{t-d})$ is the function constructs the threshold variable; $d$ denotes delay parameter with positive integer that establishes the time delay at which the threshold function $u(X_{t-d})$ significantly influences the structure of $Y_{t}$; $r\in[\text{min }u(X_{t-d}),\text{max }u(X_{t-d})]$ denotes threshold value; $p_{1}$ and $p_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $Y_{t}$, respectively; $q_{1}$ and $q_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $X_{t}$, respectively; $\{\phi_{0}^{(1)};\phi_{i}^{(1)}|i=1,2,\ldots,p_{1};\theta_{\ell}^{(1)}|\ell=1,2,\ldots,q_{1}\}$ are parameters for $1^{th}$ regime; and $\{\phi_{0}^{(2)};\phi_{i}^{(2)}|i=1,2,\ldots,p_{2};\theta_{\ell}^{(2)}|\ell=1,2,\ldots,q_{2}\}$ are parameters for $2^{nd}$ regime; $\{\varepsilon_{t}\}$ is a sequence of martingale differences satisfying ${\mathbb{E}}(\varepsilon_{t}|F_{t-1})=0$, ${sup}_{t}{\mathbb{E}}(|\varepsilon_{t}|^{\delta}|F_{t-1})<\infty$ $almost$ $surely$ for some $\delta>2$, where $F_{t-1}$ is the $\sigma$-field generated by $\{\varepsilon_{t-i}|i=1,2\ldots\}$. The unknown set of parameters for the model in Equation (4) can be shown as follows:
$\Omega=\left\{\Phi,\Theta\right\},\quad\Phi=\left\{\phi_{0}^{(1)},\phi_{i:p_{1}}^{(1)},\phi_{0}^{(2)},\phi_{i:p_{2}}^{(2)}\right\},\quad\Theta=\left\{\theta_{\ell:q_{1}}^{(1)},\theta_{\ell:q_{2}}^{(2)}\right\}.$ (5)
The hyperparameters of the model can be collected as;
$\Gamma=\{p_{1},p_{2},q_{1},q_{2},d,r\in[\text{min }u(X_{t-d}),\text{max }u(X_{t-d})]\}.$
The parameters of the model in (4) can be estimated by LS method conditional on hyperparameters similar to minimization of the objective function shown in (2) and (3).
In order to discover TAR model-based causative interactions in a network $N$ consisting of multiple sources of information, called nodes, it is helpful to specify the following definitions regarding (4) and (5).
Definition 1
[…]
(X is threshold-connected to Y if X exerts a threshold nonlinearity effect on Y)
The first definition implies that the presence of a threshold non-linearity in the TAR model (4) would be sufficient enough to show that $X_{t}$ has a threshold conditional connectivity with $Y_{t}$. Note that, as per this definition, the set of parameters $\Theta$ is considered an empty set. In other words, even if the past of $X_{t}$ does not have a significant predictive effect, which would be provided by $\Theta$, the threshold functional of $X_{t-d}$ has a particular influence on the structure of $Y_{t}$ in the sense that it controls which regime $Y_{t}$ will follow. To explore the threshold-connectivity (TC) stated in definition 1, we utilize the threshold non-linearity test scheme suggested by Hansen (1996, 2000).
The model given in (4) can also be rewritten to illustrate Hansen’s test in a simpler way. Let $W_{t}=(1\text{ }Y_{t-1}\text{ }\ldots\text{ }Y_{t-p}\text{ }X_{t-1}\text{ }\ldots\text{ }X_{t-q})^{\prime}$ be a $m\times 1$ vector where $p_{1}=p_{2}=p$; $q_{1}=q_{2}=q\text{ }\text{ }\text{ }\text{ }$ and $m=p+q+1$; $\pi_{1}=(\phi_{0}^{(1)}\text{ }\phi_{1}^{(1)}\text{ }\ldots\text{ }\phi_{p}^{(1)}\text{ }\theta_{1}^{(1)}\text{ }\ldots\text{ }\theta_{q}^{(1)})^{\prime}$ and $\pi_{2}=(\phi_{0}^{(2)}\text{ }\phi_{1}^{(2)}\text{ }\ldots\text{ }\phi_{p}^{(2)}\text{ }\theta_{1}^{(2)}\text{ }\ldots\text{ }\theta_{q}^{(2)})^{\prime}$ then the 2 regime TAR model in (4) takes the form;
M2: $Y_{t}=\pi_{1}^{\prime}W_{t}I(u(X_{t-d})\leq r)+\pi_{2}^{\prime}W_{t}I(u(X_{t-d})>r)+\varepsilon_{t}.\phantom{AAAAAA}$ (6)
In the absence of threshold nonlinearity, $\pi_{1}=\pi_{2}$ (i.e., $r=r_{0}$) the linear autoregressive distributed lag (ADL) model can be expressed as follows:
M1: $Y_{t}=\pi_{1}^{\prime}W_{t}+\varepsilon_{t}.\phantom{AAAAAAAAAAAAAAAAAAAAAAAAAAAA}$ (7)
The unknown parameters of the models represented in (6) and (7) can be shown as $\Pi_{\text{\tiny M2}}=[\pi_{1},\text{ }\pi_{2},\text{ }r,\text{ }d]$ and $\Pi_{\text{\tiny M1}}=[\pi_{1}]$, respectively. Under the assumption ${\mathbb{E}}(\varepsilon_{t}|F_{t-1})=0$ the appropriate estimation method is LS method for both. The LS estimator of $\widehat{\Pi}_{\text{\tiny M1}}$ and $\widehat{\Pi}_{\text{\tiny M2}}$, which is conditional on fixed hyperparameters $\{r,d\}$, can be found by the minimization of following objective functions;
$\widehat{\Pi}_{\text{\tiny M1}}=\operatorname*{\arg\min}_{\Pi_{\text{M1}}}\sum_{t=1}^{T}\big{[}Y_{t}-\pi_{1}^{\prime}W_{t}\big{]}^{2}$ (8)
$\widehat{\Pi}_{\text{\tiny M2}}=\operatorname*{\arg\min}_{\Pi_{\text{M2}}}\sum_{t=1}^{T}\big{[}Y_{t}-\pi_{1}^{\prime}W_{t}I(u(X_{t-d})\leq r)-\pi_{2}^{\prime}W_{t}I(u(X_{t-d})>r)\big{]}^{2}$ (9)
Models given in (6) and (7), which is being the most restricted, have a nested structure. Thereby, comparing them using a model selection criterion is equivalent to a likelihood ratio test (Hansen, 1999).
Let $S_{\text{\tiny M1}}={\hat{\varepsilon}_{\text{\tiny M1}}}^{\prime}\times{\hat{\varepsilon}_{\text{\tiny M1}}}^{\text{}}$ and $S_{\text{\tiny M2}}=\hat{\varepsilon}^{\prime}_{\text{\tiny M2}}\times{\hat{\varepsilon}_{\text{\tiny M2}}}^{\text{}}$ are denoting sum of squared residuals obtained by LS estimation in (8) and (9), respectively; and let $r_{0}$ is denoting the true value of the threshold value $r$; then testing the existence of threshold non-linearity can be accomplished by the testing the hypothesis of $\text{H}_{0}:r=r_{0}$ considering the following likelihood ratio (LR) test statistic;
$LR(r)=T\bigg{(}\frac{S_{\text{\tiny M1}}-S_{\text{\tiny M2}}(\hat{r})}{S_{\text{\tiny M2}}(\hat{r})}\bigg{)}.$ (10)
The test of the hypothesis of $\text{H}_{0}$ (i.e., M1 against M2 or linear vs. non-linear model) is to reject for large values of $LR(r_{0})$. This is the LR test under the auxiliary assumption that $\varepsilon_{t}$ are independent $N(0,\sigma^{2})$. It also corresponds to the conventional F (or Wald) and the Lagrange multiplier (or score) tests. However, the asymptotic distribution of the $LR$ does not follow a chi-square distribution since the nuisance parameter $r$ is not identified and does not enter the model under the null hypothesis of no threshold effect. To overcome this barrier, Hansen (1996) demonstrates that the asymptotic distribution and thus $p$-values can be approximated by a residual bootstrap simulation. The test procedure also provides a heteroskedasticity-consistent Lagrange multiplier test that adjusts for heteroskedastic error terms.
The asymptotic distribution of the threshold nonlinearity test and some of the major theoretical problems of inference for TAR model estimation are comprehensively studied by Hansen (1996, 1997, 1999, 2000). Through testing the existence of threshold non-linearity, we aim to discern the connectivity between channels, shedding light on the directed connectivity within the brain.
Definition 2
[…]
(X threshold Granger causes Y if X has both threshold nonlinearity effect on Y and also having predictive power on Y)
With our second definition, we specify the threshold Granger causality (TGC) by incorporating the impact of the threshold node’s predictive ability in the TAR model. Once the threshold nonlinearity is verified, the form of the causal interactions between the two nodes of a network can be explored and represented by terms of the TAR model. In addition to the previously defined threshold-connectivity, in this second definition, the threshold node $X_{t-d}$, whose threshold connectivity to $Y_{t}$ is determined by the threshold non-linearity test, is included as a predictor of the current state of $Y_{t}$. Hence, the identification of TGC would require the presence of statistically significant coefficients for any of the time-lagged terms of the $X_{t}$. Given the possibility of multiple valid representations of causal interactions, it is necessary to establish a model selection criterion that avoids over-parameterization for a well-refined representation of the causal connections. To this end, we suggest employing the heteroscedasticity-robust Wald statistics provided by Hansen (1997) be used as a term to be penalized in model selection criteria. The Wald test for causality can be structured to test the predictive capacity of the threshold variable included as explanatory covariates in the model. It should be noted that Li (2006) was the first to mention the TGC nomenclature and made use of the Wald test, proposed and studied in detail by Hansen (1996, 1997) for the TAR model specification (i.e., testing the restricted vs unrestricted models). In his model, Li (2006) utilized the dependent variable itself as the threshold variable. Therefore, it is worth mentioning that whilst we adopt the common TGC nomenclature, our suggested methodology diverges in terms of the causal connectivity framework.
Considering the model given in (4) the hypothesis to be tested for TGC can be expressed as follows:
$\text{H}_{0}:\theta_{1}^{(1)}=\theta_{1}^{(2)}=\ldots=\theta_{q_{1}}^{(1)}=\theta_{q_{2}}^{(2)}=0,$ (11)
where the alternative hypothesis discerns the TGC by stating that at least one coefficient is nonzero.
If the threshold value $r$ splitting the sample and delay $d$ are known, the pointwise heteroscedasticity-consistent Wald statistic can be shown as follows:
$\Delta_{T}(r,d)=\bigg{(}\text{C}\hat{\Omega}(r,d)\bigg{)}^{\prime}\bigg{[}\text{C}\bigg{(}M(r,d)^{-1}V(r,d)M(r,d)^{-1}\bigg{)}\text{C}^{\prime}\bigg{]}^{-1}\bigg{(}\text{C}\hat{\Omega}(r,d)\bigg{)},$ (12)
where C is the selector matrix; $M(r,d)=\sum_{t=1}^{T}W_{t}(r,d)W_{t}(r,d)^{\prime}$;
$V(r,d)=\sum_{t=1}^{T}W_{t}(r,d)W_{t}(r,d)^{\prime}\hat{\varepsilon_{t}}^{2}$; $W_{t}(r,d)=\bigg{(}W_{t}^{\prime}I(u(X_{t-d})\leq r)\text{ }\text{ }W_{t}^{\prime}I(u(X_{t-d})>r)\bigg{)}^{\prime}$, then according Hansen (1996); Davies (1977, 1987) the appropriate testing of $\text{H}_{0}$ can be done by using following LR statistic
$\Delta_{T}=\sup_{(r,d)\in\Gamma}\Delta_{T}(r,d).$ (13)
In contrast to the test statistic given in (10), this testing treats $(r,d)$ as fixed in both the null and alternative hypotheses, which makes it standard testing procedure, and the asymptotic distribution of $\Delta_{T}$ is $\chi^{2}_{\rho(C)}$, where ${\rho(C)}$ is the number of restrictions. The test rejects the null of the restricted model for the large values of $\Delta_{T}$. It should be noted that the restricted model tested via the hypothesis given in (11) is not the sole hypothesis implying TGC. All alternatives to the restricted models stated in the null hypothesis with respect to the $\Theta$ would imply TGC. In fact, the restricted models to be tested that indicate TGC should be investigated through a grid search since the statistic $\Delta_{T}$ varies over hyperparameters $p_{1},p_{2},q_{1},q_{2}$. Further, $\Delta_{T}$ provides the deviance between the restricted vs. unrestricted model, which makes it a proper tool to employ in model selection criteria for selecting a simpler one among statistically valid models. Thereby, following the confirmation of threshold nonlinearity, to ascertain a parsimonious manifestation of the causal interaction between $X$ and $Y$, we propose a model selection criterion, which is a modified version of the Schwarz criterion (SC),
${mSC}_{T}=\Delta_{T}-{2log(T)K},$ (14)
with $K$ is the number of parameters in the model.
The model selection criterion based estimation of TGC can be defined as the maximization of the objective function given in (14),
$\max_{\Gamma}\text{ }{mSC}_{T}.$ (15)
The maximization of (14) yields estimates of parameters $\hat{\Omega}$ and $\hat{\Gamma}$ as a by-product. The use of a method of model selection for inference with a criterion analogous to that in (14) and related remarks have been given by Raftery (1995), Kass and Raftery (1995), Liu et al. (1997), and Gonzalo and Pitarakis (2002).
The whole procedure referred to as TAR4C can be simplified as follows:
Network : $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
Scope : $\text{Analyze}\text{ }{\text{TC}}_{X\to Y}\text{ and}\text{ }{\text{TGC}}_{X\to Y}\text{ }\text{ }$
Input : $X,Y\in N$ (nodes of interest in a network, $N$)
Parameters : $\Omega,\Gamma$
Data : $\{({X}_{t},{Y}_{t})\}_{t=1}^{T}$ the set of obs. of two nodes from the network $N$;
$u({X}_{t-d})$ threshold functional
Decision : (a) ${\text{TC}}_{X\to Y}$ (b) ${\text{TGC}}_{X\to Y}$ (c) ${\text{TC}}_{X\not\to Y}$
1
2Step 1. Neutralize interference from other nodes:
  1. Step 1.1.
    $Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|H(Z_{1,t},\ldots,Z_{N-2,t})\big{)}$
  2. Step 1.2.
    $X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|H(Z_{1,t},\ldots,Z_{N-2,t})\big{)}$
  3. $H(.)$ denotes for spectral domain dynamic PCA.
  • 3Step 2.
    Testing threshold effect:
    1. Step 2.1.
      Specify the $u(X_{t-d})$ of the node $X$ whose causative effect on $Y$ is of interest.
    2. Step 2.2.
      Regarding the model in Equation (6), test the following hypothesis of linearity.
      $\text{H}_{0}:\pi_{1}=\pi_{2}$
    3. Step 2.3.
      Any significant evidence against the hypothesis of linearity indicates ${\text{TC}}_{X\to Y}$. If the decision is in favor of threshold non-linearity, proceed to Step 3. Otherwise, or for different causality assessments, repeat Step 2 according to different threshold functionals.
    4. Step 2.4.
      If all else fails, the decision ${\text{TC}}_{X\not\to Y}$ can be imposed.
  • 4Step 3.
    Identify causal representation by TAR:
  • 5
    1. Step 3.1.
      $\max_{\Gamma}\text{ }{mSC}_{T}$
    2. Step 3.2.
      Obtain $\hat{\Omega}$ and $\hat{\Gamma}$ as a by-product of the maximization.
    3. Step 3.3.
      ${\text{TGC}}_{X\to Y}$ if $\hat{\Theta}\neq\{\emptyset\}$
    6
    Algorithm 1 TAR4C$(X,Y)$
    2.4 On the role of TAR models for exploring causal connectivity in a network of multiple nodes
    Introducing TAR models into causality analysis for networks extends their applicability to interconnected systems. This approach offers researchers new perspectives to explore complex interactions, highlighting threshold nonlinear causality in dynamic environments. Unlike conventional Granger causality tests, which rely on linear assumptions, the use of TAR models allows for the detection of nonlinear causal relationships between variables. This is particularly important as many real-world relationships exhibit nonlinearities, and detecting such relationships enhances the scope of causality analysis. The inclusion of a threshold nonlinearity test helps identify points where causal connections between nodes change qualitatively, enhancing interpretability in dynamic networks like EEG. Identifying regime-switching dynamics adds a layer of interpretability about the nature of causal impact in a dynamic network such as the EEG setting and can offer insights into the conditions under which, for example, the node $X$ influences the node $Y$.
    TAR models accommodate various regime-specific dynamics, offering a more refined representation of causal interactions under different conditions, where linear models may fall short. This flexibility allows experts to tailor the model to the data, improving representation of the underlying causal mechanisms. However, TAR models can be more complex, potentially leading to over-fitting or over-parameterization, especially with limited data. Estimating TAR models, particularly identifying thresholds, can also be computationally intensive, posing challenges for large networks or datasets.
    3 Application to Real Data
    3.1 Dataset and the experimental procedure
    In this section, we exemplify the utility of the proposed approach for identifying connectivities and causal interactions in a complex time series network by implementing it on a real-world EEG dataset. The dataset used in the analysis originated from an experiment that entailed a series of motor/imagery tasks, providing rich information on the temporal dynamics of brain activity. The whole experiment consisted of EEG recordings from 109 volunteers using a 64-channel setup and publicly available at PhysioNet.org (Schalk et al., 2004). The experimental protocol comprised 14 experimental runs for each subject, including two baseline runs (one with eyes open and one with eyes closed) and twelve task-oriented runs. These tasks required participants to either physically perform or imagine movements in response to visual cues on a screen, such as opening and closing fists or feet, depending on the position of a target. The tasks are divided into four categories: physical movement of opening and closing of the left-hand or right-hand fist, imaginary movement of opening and closing the left-hand or right-hand fist, physical movement of opening and closing of both fists’ or both feet’, and imaginary movement of opening and closing of both fists’ or both feet’. EEG data were sampled at 160HZ, along with an annotation channel indicating rest or the onset of motion (real or imagined) for specific movements. The EEG montage followed the international 10-10 system, omitting certain electrodes (i.e., Nz, F9, F10, FT9, FT10, A1, A2, TP9, TP10, P9, and P10).
    All available data underwent the pre-processing phase to filter out artifacts (i.e., non-brain related activities) only, such as eye blinks and muscle-related movements, thereby avoiding interference that would damage any temporal structure of the recordings. The pre-processing procedure consists of the following steps: Initially, the recordings are subjected to a high-pass filter at 0.5Hz to eliminate low-frequency noise. This is followed by the removal of AC power-related line noises at 60Hz. Next, Independent Component Analysis (ICA), facilitated by MNE-python routines (Gramfort et al., 2013; Larson et al., 2024), is implemented to isolate and exclude the source of artifact components. Finally, artifacts associated with eye blinking and muscle movements are identified and removed, ensuring the recordings are free from these specific types of interference.
    Despite all subjects’ recordings undergoing the specified standard pre-processing procedure, some structural problems, such as sawtooth trace, excessive blinking, broken channels, unbalanced epoching, etc., could not be resolved and were identified in some subjects’ records. Therefore, we limited the scope of our analysis to 35 Right-handed subjects (17 Females and 18 Males) whose recordings were found to be qualified. Further, to more appropriately demonstrate how the proposed method can be used to dissect causal connections between pairwise nodes within any network of time series, we focused on the associations among four nodes (i.e., channels), with two (i.e., AF3, AF4) positioned over the pre-frontal cortex and two (i.e., C3, C4) placed over the motor cortex illustrated in Fig. 1. These channels were chosen because they correspond to regions of the brain (i.e., prefrontal and motor cortices) that are likely to be active during voluntary movement and imagery.
    Figure 1: Channels of interest colored on a sample of 64-channel EEG placement.
    The connectivities between these channels were examined in 35 subjects’ data pertaining to both motor movement (MM) and motor imagery (MI) tasks. The subjects performed a series of opening and closing (OC) gestures, including left-hand fist (LH), right-hand fist (RH), both fists’ (BH), and both feet’ (BF), for both the motor movement (MM) and motor imagery (MI) tasks. The aim was to compare the brain activity associated with these tasks and gestures by identifying the patterns of connectivity and causal interactions amongst the selected channels. The connectivities explored can be outlined in Table 1.
    Table 1: Outline for illustrating causal connections during different brain activities.
    Motor Movement Motor Imagery
    LH [Uncaptioned image] [Uncaptioned image]
    RH [Uncaptioned image] [Uncaptioned image]
    BH [Uncaptioned image] [Uncaptioned image]
    BF [Uncaptioned image] [Uncaptioned image]
    *The black arrows indicate the pairwise causal interactions explored.
    3.2 Implementation of the proposed approach
    Each subject repeated the relevant task (i.e., task1: OC of LH/RH, task2: imagery OC of LH/RH, task3: OC of BH/BF, task4: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor movement/imagery motion (i.e., LH, RH, BH, BF) approximately 23 times in total.
    Each epoch consists of about 4 seconds, that is, 640 observations. Initially, the relevant epochs were extracted from the dataset with the help of annotation information. For each epoch, 8 different causality interactions were tested pairwisely. For each task, 23 test results were combined using the method suggested by Becker (1994), and a decision was made on the presence of Threshold Non-linearity in the relevant task for the subject.
    The proposed method for inferring connectivity for just one pair of channels has been represented in Chapter 2 by the procedure named TAR4C. In the application to the whole dataset, we first focused on 2 index metrics in order to provide connectivity schemes covering the analysis results of all subjects considered.
    3.2.1 Threshold Connectivity Index
    The Threshold Connectivity Index (TCI) quantifies, considering all the subjects, the presence of connectivity between two EEG channels based on the detection of threshold nonlinearity in their interaction. This index reflects the overall connectivity between the channels under specific conditions, such as during motor imagery or motor movement tasks.
    The TCI is calculated as follows:
    $\text{TCI}_{XY}=\left(\frac{1}{12\times N}\sum_{d=1}^{12}\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)\cdot n^{(d)}\right)\times 100,$
    where $n^{(d)}$ is the number of subjects out of the total $N$ who reject the linearity test at delay $d$. This indicates the number of subjects exhibiting threshold nonlinearity at the specified delay. $\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)$ is an indicator function. It equals 1 when $n^{(d)}\geq 0.7\times N$, meaning that the number of subjects rejecting linearity at delay $d$ equals or exceeds 70% of the total subjects. If this condition is not met, the indicator function equals 0. This ensures that only delays where a significant proportion of subjects (at least 70%) reject linearity are included in the TCI calculation. The resulting TCI value reflects the proportion of subjects showing significant threshold connectivity between channels $X$ and $Y$ and is expressed as a percentage. In practical terms, a higher TCI value suggests stronger threshold-based connectivity between the two channels, which is visualized in network diagrams with thicker lines representing stronger connections. The analysis procedure for calculating TCI involves the following steps:
    Table 2: Outline of the implementation for TCI calculation.
    Motor Movement / Motor Imagery
[…]
    3.2.2 Threshold Granger Causality Index
    The Threshold Granger Causality Index (TGCI) extends the TCI by incorporating an additional step that tests the predictive power of the threshold variable. After establishing the presence of threshold nonlinearity, the TGCI assesses whether the past values of the threshold variable significantly predict the future values of the dependent variable using the Wald test. This index, therefore, provides a measure of directional causal influence between EEG channels.
    The TGCI is calculated as follows:
    $\text{TGCI}_{XY}=\left(\frac{1}{12\times N}\sum_{d=1}^{12}\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)\cdot\text{Wald}^{(d)}\right)\times 100,$
    where $\text{Wald}^{(d)}$ is the number of subjects out of the total $N$ who pass the Wald test in favor of the predictive power of the threshold variable $X$ at delay $d$. This index quantifies the strength of the causal influence from $X$ to $Y$.
    The TGCI reflects the strength and directionality of the causal relationship between the channels and is used to visualize the directional flow of information in network diagrams. A higher TGCI value indicates a stronger causal influence from the threshold variable to the dependent variable.
    The analysis procedure for calculating TGCI builds on the steps used for TCI and includes the following additional steps:
    Table 3: Outline of the implementation for TGCI calculation.
    Motor Movement / Motor Imagery
    Gesture to Perform Let $X$ and $Y$ represent two nodes in the EEG network. Compare whether $X$ is Threshold G - Cause $Y$ using the proposed approach given in Section 2.3. 1. Follow the steps as outlined in the TCI calculation procedure. 2. For each delay value $d$ where threshold nonlinearity is detected, apply the Wald test to assess the predictive power of the threshold variable $X_{t-d}$. 3. Combine the Wald test results over approximately 23 epochs for each subject. 4. Calculate the TGCI by averaging the Wald test results across all subjects and normalizing by the number of delays and subjects. 5. Visualize the TGCI results to identify the direction and strength of causal influence between EEG channels.
    3.3 Connectivity at Motor Imagery
    MI left-hand fist gesture (MI-LH): For the MI-LH gesture, the TCI results (Table 4) indicate moderate connectivity between the channels. The connectivity from AF3 to C3 (TCI = 7.14, red arrow) and from AF4 to C4 (TCI = 7.86, red arrow) is slightly higher
    TCI at MI TGCI at MI
    LH
    [Uncaptioned image] [Uncaptioned image]
    RH
    [Uncaptioned image] [Uncaptioned image]
    BHs
    [Uncaptioned image] [Uncaptioned image]
    BF
    [Uncaptioned image] [Uncaptioned image]
    Table 4: Connectivity Nets at Motor Imagery
    than the reverse connections (C3 to AF3, TCI = 6.67, blue arrow; C4 to AF4, TCI = 7.62, blue arrow). For AF4, the TCI values show 6.19 (red arrow) from AF4 to C3 and 6.43 (blue arrow) from C3 to AF4, indicating nearly equal connectivity in both directions. These results suggest somewhat balanced connectivity between these channels, with a slight prominence of the motor cortex over the prefrontal cortex in the C4 to AF3 (TCI = 14.29, blue arrow).
    The TGCI values for the MI-LH gesture (Table 4) show a more complex pattern. While the TCI indicated moderate connectivity, the TGCI values reveal that the causal influence from C4 to AF3 (TGCI = 12.62, blue arrow) is stronger than its reverse direction from AF3 to C4 (TGCI = 5.95, red arrow), and the causal influence from AF4 to C4 is notably high (TGCI = 25.24, red arrow) than its opposing direction from C4 to AF4 (TGCI = 13.1, red arrow). These indicate a significant role of AF4 and C4 in driving motor activity during left-hand fist motor imagery.
    MI right-hand fist gesture (MI-RH): The MI-RH gesture results exhibit stronger overall connectivity, as seen in the TCI values(Table 4). The connection from AF3 to C3 has a TCI of 13.10 (red arrow), while the reverse direction, C3 to AF3, has a TCI of 5.95 (blue arrow), indicating a stronger connectivity from AF3 over C3. The connection from AF3 to C4 has a TCI of 14.29 (red arrow), with the reverse, C4 to AF3, showing 6.67 (blue arrow), again highlighting stronger prefrontal influence. The connections involving AF4 also show similar patterns, with TCI values of 13.33 (red arrow) from AF4 to C3 and 6.19 (blue arrow) in the reverse direction, and 13.57 (red arrow) from AF4 to C4 with 12.62 (blue arrow) in the reverse.
    The TGCI values for MI-RH (Table 4) also reflect strong directional connectivity, particularly from the channels over the prefrontal to the motor cortices. AF3 to C4 and AF4 to C4 both have TGCI values of 18.33 (red arrows), indicating significant causal influence from the channels over the prefrontal cortex to the motor regions. The reverse connections, C4 to AF3 and C4 to AF4, both have lower TGCI values of 6.19 (blue arrows), suggesting that while there is some feedback, it is much weaker compared to the prefrontal influence.
    MI both hands’ fists gesture (MI-BHs) For the MI-BHs gesture, the TCI results (Table 4) show balanced and strong connectivity between the channels over the prefrontal and the motor cortices. The connection from AF3 to C3 has a TCI of 7.38 (red arrow), with the reverse direction showing 6.90 (blue arrow). The connection from AF3 to C4 has a TCI of 14.05 (red arrow), while C4 to AF3 is 13.57 (blue arrow), indicating nearly equal connectivity in both directions. The connectivity involving AF4 is similarly strong and balanced, with TCI values of 14.05 (red arrow) from AF4 to C3 and 12.62 (blue arrow) in the reverse direction, and 13.81 (red arrow) from AF4 to C4 with 7.86 (blue arrow) in the reverse.
    The TGCI results for MI-BH reveal significant contralateral causal influence, particularly between C3 and AF4 (TGCI = 17.86, blue arrow) and AF3 to C4 (TGCI = 12.14, red arrow). The connection from C3 to AF4 and having no flow in the opposite direction of itself indicate a strong causal influence from the channel over the left motor cortex to the right prefrontal cortex.
    MI both feet gesture (MI-BF) The MI-BF gesture (Table 4) shows the highest TCI values, particularly for the connections between AF3 and C4. The connection from AF3 to C4 has a TCI of 21.19 (red arrow), with the reverse, C4 to AF3, showing 20.95 (blue arrow), indicating exceptionally strong and nearly equal connectivity in both directions. The connections from AF4 to C3 and C4 also show high TCI values, with 13.10 (red arrow) from AF4 to C3 and 13.81 (blue arrow) in the reverse direction, and 14.76 (red arrow) from AF4 to C4 with 14.52 (blue arrow) in the reverse, indicating strong and balanced connectivity.
    The TGCI results further emphasize this strong connectivity, with particularly high values for the feedback from C3 to AF4 (TGCI = 24.05, blue arrow) and from C4 to AF4 (TGCI = 18.10, blue arrow). These findings indicate that during both feet’ motor imagery, there is a continuous exchange of information between the channels over motor and prefrontal cortices, with the motor regions having a significant causal influence over the prefrontal areas. This strong bidirectional connectivity is crucial for the precise coordination required for bilateral foot movements.
    Comparative Analysis Between gestures:
    When comparing the MI-LH and MI-RH gestures, the RH gesture exhibits stronger overall connectivity. The TCI values for the RH gesture are higher across nearly all channel pairs compared to the LH gesture. For instance, the TCI from AF3 to C3 is 13.10 (red arrow) in the RH task compared to 7.14 (red arrow) in the LH task. Similarly, the connectivity from AF3 to C4 is stronger in the RH task (TCI = 14.29, red arrow) than in the LH task (TCI = 6.90, red arrow). This suggests that the RH motor imagery, which typically involves the dominant hand in most subjects, engages more robust neural connectivity, particularly from the channels over the prefrontal cortex to the motor cortices.
    The TGCI results further emphasize this difference. The MI-RH shows strong directional causal influence from AF3 and AF4 to C4, with TGCI values of 18.33 (red arrows) for both connections, indicating that the channels over the prefrontal cortex hold significant causal control over the motor cortex during this task. In contrast, the MI-LH shows weaker causal connectivity, with the most significant TGCI being 25.24 (red arrow) from AF4 to C4, but less pronounced overall compared to the RH task. This difference likely reflects the increased cognitive demand and neural resources required for controlling the dominant hand during motor imagery.
    The MI-BF gesture shows the highest TCI values, particularly in the interactions between AF3 and C4, and between C4 and AF3. This suggests that the motor imagery of both feet’ gesture involves stronger connectivity compared to other gestures, likely due to the increased complexity of coordinating bilateral foot movements.
    Across all gestures, a few general trends can be observed. First, bilateral tasks (BHs and BF) generally involve stronger and more complex channel connectivity compared to unilateral tasks (LH and RH). This is reflected in both the TCI and TGCI values, where bilateral tasks show higher indices, indicating more extensive communication and coordination between channels. This increased connectivity likely reflects the higher cognitive demands associated with coordinating movements across both sides of the body.
    Second, channels over the prefrontal cortex appear to play a more noticeable role in controlling channels over motor regions during right-hand fist tasks, as evidenced by the higher TCI and TGCI values for connections from AF3 and AF4 to C4 in the right-hand fist gesture. This prominence is less pronounced in the left-hand fist gesture, where connectivity and causal influences are generally weaker.
    Finally, the bidirectional connectivity observed in TGCI results, particularly during both fists’ and both feet’ gestures, highlights the significance of feedback between the channels over motor and over prefrontal regions during complex motor imagery tasks.
    3.4 Connectivity at Motor Movement
    MM left-hand fist gesture (MM-LH): For the MM-LH, the TCI results (Table 5) indicate pronounced connectivity from the channels over motor cortex back to the channels over prefrontal cortex, as evidenced by the high TCI value of 18.81 from C3 to AF3 (blue arrow). This is significantly higher than the connectivity from AF3 to C3 (TCI = 7.14, red arrow), suggesting that during actual left-hand fist movement, channel C3 maintains stronger connectivity on channel AF3 rather than the reverse. Similar patterns are observed between AF3 and C4, with a TCI of 13.81 (red arrow) from AF3 to C4 and a slightly lower TCI of 12.14 (blue arrow) from C4 to AF3. The connectivity between AF4 and C3/C4 is more balanced, with nearly equal TCI values in both directions.
    The TGCI results for MM-LH further highlight this motor-to-prefrontal influence, with a strikingly high TGCI of 25.00 (blue arrow) from C3 to AF3, indicating a strong causal influence. The causal influence from the prefrontal cortex to the motor regions is relatively weaker, as shown by the TGCI values of 5.95 (red arrow) from AF4 to C3. This suggests that during left-hand fist movement, the channels over the motor cortex are highly connected to and causally influential over the channels AF3 and AF4, likely reflecting the real-time processing required for executing the movement.
    TCI at MM TGCI at MM
    Left Fist
    [Uncaptioned image] [Uncaptioned image]
    Right Fist
    [Uncaptioned image] [Uncaptioned image]
    Both Fists
    [Uncaptioned image] [Uncaptioned image]
    Both Feet
    [Uncaptioned image] [Uncaptioned image]
    Table 5: Connectivity Nets at Motor Movement
    MM right-hand fist gesture (MM-RH): The MM-RH gesture exhibits more balanced connectivity between the channels. The TCI values (Table 5) show significant connectivity from AF3 and AF4 to both C3 and C4, with TCI values ranging from 7.14 to 14.29 (red arrows). The reverse connections, particularly from C3 to AF3 (TCI = 12.86, blue arrow) and from C3 to AF4 (TCI = 12.86, blue arrow), indicate a bidirectional connectivity.
    TGCI results for MM-RH reveal strong causal connectivity from AF3 to C4, with TGCI values of 18.33 (red arrow). This suggests that during right-hand fist movement, channels AF3 and AF4 play a role in driving motor activity. However, the reverse causal influence from C3 and C4 to Af3 and AF4 is less pronounced, as indicated by the lower TGCI values, such as 5.95 (blue arrow) from C3 to AF3. This pattern highlights the prominence of the prefrontal cortex in controlling motor functions during right-hand fist movement, reflecting the cognitive demand associated with executing movements of the dominant hand.
    MM both hand’s fists gesture (MM-BHs): For both fists’ gesture, the TCI results (Table 5) show relatively proportional connectivity across the prefrontal and motor cortices. The TCI values from AF4 to C3 is 14.29 (red arrow), and from AF3 to C4, it is 13.33 (red arrow), suggesting substantial connectivity over C3 and C4.
    However, the TGCI results for the both fists’ gesture seem to oppose this finding, with significant causal influence observed from C4 to AF4 (TGCI = 12.14, blue arrow) and from C3 to AF3 (TGCI = 11.90, blue arrow). The contrast in terms of TCI and TGCI during both fists’ movement suggests there is a continuous exchange of information between the channels over prefrontal and motor regions, which might be related to coordinating bilateral motor activity. The bidirectional causal connectivity suggests that both hemispheres are actively engaged, with the prefrontal cortex contributing to motor planning and the motor cortex providing feedback to adjust and refine the movements.
    MM both feet gesture (MM-BF): The MM both feet’ gesture shows the highest TCI values (Table 5) among all gestures, particularly for the connections from C3 to AF3, with a TCI of 19.29 (blue arrow), and from AF3 to C4, with a TCI of 14.29 (red arrow). These results suggest that the MM both feet’ task involves strong connectivity between the prefrontal and motor cortices, especially between the left prefrontal and both motor areas.
    The TGCI results for MM-BF support this interpretation, showing very high TGCI values, especially from C3 to AF3 (TGCI = 31.67, blue arrow). This indicates a principal causal influence of the C3 over the AF3 during MM-BF, highlighting the significant feedback mechanisms necessary for executing and adjusting bilateral feet movement. The causal influence from AF3 to C3 and C4 is also strong, particularly from AF3 to C4 (TGCI = 19.29, red arrow), reflecting the prefrontal cortex’s role in controlling these complex movements.
    Comparative Analysis Between gestures: Across all gestures, several key trends emerge. First, the overall connectivity strength, as measured by TCI, tends to be higher for both fists’ and both feet’ gestures than the left-hand fist and right-hand fist gestures. This suggests that bilateral motor tasks require more extensive neural coordination, with stronger interactions between the prefrontal and motor cortices.
    Second, the TGCI results indicate that during unilateral tasks (left-hand fist and right-hand fist), the motor cortex wields a more prominent causal influence over the prefrontal cortex, especially in the left-hand fist gesture. This pattern likely reflects the need for real-time feedback and adjustments during the execution of these movements, particularly when performing tasks with the non-dominant hand (left-hand fist).
    Finally, the bidirectional connectivity observed in TGCI results during both fists’ and both feet’ gestures underlines the importance of continuous feedback and coordination between the channels over prefrontal and motor regions during these complex bilateral tasks. This feedback mechanism suggests that both hemispheres are involved in synchronizing movements, which appears to be necessary for the successful execution of coordinated bilateral actions.
    The connectivity patterns across the different gestures highlight the varying demands on neural coordination and causal influence depending on the specific motor task. Bilateral movements (both fists’ and both feet’) involve stronger and more complex connectivity, reflecting the higher cognitive and motor demands associated with coordinating movements across both sides of the body. Unilateral movements, particularly those involving the non-dominant hand, require significant motor-to-prefrontal feedback, emphasizing the importance of real-time adjustments during these tasks.
    3.5 Comparative Analysis of Threshold Granger Causality: Motor Imagery vs Motor Movement
    We have thus far presented and discussed the connectivity between the channels separately for MI and MM tasks using TCI and TGCI values. The figures presented in Table 4 and Table 5 visualize the corresponding results. Furthermore, we conducted an additional analysis to assess the differences in causal connectivities within and between MI and MM tasks, employing a permutational Hotelling’s $T^{2}$ test (Curran and Hersh (2021)) on data from 35 subjects. In these tests, the connectivity of each pair within MI and MM was compared with the connectivity of the same pair in the opposite direction. To illustrate, the TGC sampling distributions of MI/MM at $X$ to $Y$ were tested against the TGC sampling distributions of MI/MM at $Y$ to $X$. The figures given in Table 7 display the significant differences in connectivities during MI and those during MM in columns 1 and 2, respectively. The differences between MI and MM were estimated by comparing the connectivity distributions in the same direction and provided by column 3. That is, the TGC sampling distributions at $X$ to $Y$ during MI were compared to the TGC sampling distributions during MM at $X$ to $Y$. Each arrow in the table signifies a statistically significant difference in directional connectivity, with the color indicating the direction of influence: red arrows represent prefrontal-to-motor differences, blue arrows indicate motor-to-prefrontal differences, green arrows signify stronger connectivity in favor of MI, purple arrows indicate stronger connectivity in favor of MM. The implemented procedure followed during the comparison is outlined in Table 6:
    Table 6: Outline of the Implementation for Comparative Analysis of TGC differences
    Gesture to Perform
    For all subject; Let $X$ and $Y$ represent two nodes in the EEG network. 1. Apply the threshold nonlinearity test using the 12th lag TAR model for each epoch and combine test results. Then, generate a 35 (i.e., number of subjects) by 1 vector for each delay $d$. (a) Retain the F-test statistics for subjects where linearity is rejected and assign 0s for others. 2. Apply the Wald test on the data of subjects who reject linearity to assess the predictive power of the threshold variable $X_{t-d}$. (a) Filter the vector of F-test statistics based on subjects who pass the Wald test. (b) Retain only those elements corresponding to subjects who pass the Wald test, forming the TGC samples for each delay $d$. 3. Compare the TGC sampling distributions for each pairwise channel interaction $X$ to $Y$ vs $Y$ to $X$ within the same task (i.e., MI or MM) using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in the directional connectivity between $X$ to $Y$ vs $Y$ to $X$ within the same task (i.e., MI or MM). (b) Visualize significant differences with arrows: Red arrows for significant prefrontal-to-motor connectivity. Blue arrows for significant motor-to-prefrontal connectivity. 4. Compare the TGC sampling distributions in the same direction (e.g., $X$ to $Y$) across MI and MM tasks using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in connectivity between MI and MM for each channel pair. (b) Visualize significant differences with arrows: Green arrows for stronger connectivity at MI compared to MM. Purple arrows for stronger connectivity at MM compared to MI.
    Table 7: Threshold Granger Causality Differences for Motor Imagery (MI), Motor Execution (MM), and MI vs MM across different motor tasks
    Differences at MI Differences at MM Differences MIvsMM
    Left Fist
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    Right Fist
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    Both Fists
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    Both Feet
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    For the left-hand fist MI task, the statistically significant differences in TGC samples given in Table 7 show bidirectional connectivities between AF4 and C4, with a red arrow from AF4 to C4 and a blue arrow from C4 to AF4. The TGCI results support this finding, indicating a strong connection from AF4 to C4 (TGCI = 25.24, red arrow) and notable feedback from C4 to AF4 (TGCI = 13.10, blue arrow). This consistency between the significant differences in TGC samples and TGCI results highlights bidirectional connectivity between these channels during MI, emphasizing the role of AF4 in influencing the C4 and the feedback mechanism. Table 7 also shows a significant connection from C4 to AF3 (blue arrow) for the MI-LH, which is indeed captured by the TGCI results, with a TGCI value of 12.62 for C4 to AF3 and 5.95 for AF3 to C4. This demonstrates that the connection from C4 to AF3 is more prominent than the reverse, indicating that C4 exerts more influence on the prefrontal region AF3 during MI-LH.
    In the MM task for left-hand fist movement, both the significant differences in TGC samples and TGCI results emphasize strong motor-to-prefrontal connectivity, particularly from C3 and C4 to AF3 (blue arrows). The TGCI values are particularly high from C3 to AF3 (TGCI = 25.00, blue arrow), highlighting the motor area channels’ significant influence on AF3 during actual movement. The red arrows from AF3 and AF4 to C4 in Table 7 may indicate the low degree of influence of the prefrontal area channels over C4 since it is not highlighted by TGCI.
    For the right-hand fist MI task, Table 7 shows significant motor-to-prefrontal influence, with blue arrows from C3 to AF3 and AF4. This dynamic does not appear to align with the TGCI outcomes, wherein C4 exerts an influence on AF3 and AF4 (TGCI = 6.19, indicated by the blue arrow). Additionally, TGCI results highlight strong prefrontal-to-motor connectivity, particularly from AF3 and AF4 to C4 (TGCI = 18.33 for both, red arrows), suggesting that during MI, AF3, and AF4 play a notable role in driving motor area channels with considerable feedback from C3 and C4. The discordance between the statistically significant differences in TGC samples and TGCI outcomes suggests that both methodologies provide complementary insights into the bidirectional connectivity essential for the intricate cognitive process of motor imagery.
    In the right-hand fist MM task, Table 7 shows significant connectivity from AF4 to C4 (red arrow), which corresponds with the TGCI results, where AF4 to C4 is also highlighted (TGCI = 6.43, red arrow). This alignment between the significant differences in TGC samples and TGCI suggests that AF4 may play a critical role in driving RH motor activity. However, the feedback from C3 to AF4 identified in Table 7 (blue arrow) does not have a direct counterpart in the TGCI results. Additionally, the TGCI results indicate a strong connection from AF3 to C4 (TGCI = 18.33, red arrow) and also a notable connection from C4 to AF3 (TGCI = 11.9, blue arrow). This finding suggests that there is significant feedback from C4 to AF3, which may play a critical role in RH motor activity. This difference indicates that while TGCI captures the influence from the prefrontal area channels to motor area channels, the significant differences in TGC samples highlight a refined bidirectional interaction, particularly from channels over the motor area to the prefrontal area, that may not be as prominent in the TGCI analysis.
    For both fists’ MI task, Table 7 highlights significant connectivity from AF3 and AF4 to C4 (red arrows), which is consistent with the TGCI results showing similar connections (TGCI = 12.14 for AF3 to C4, red arrow). This alignment between the significant differences in TGC samples and TGCI underlines the importance of AF3 and AF4 in coordinating both fists’ movement during MI. Table 7 also identifies significant motor-to-prefrontal feedback from C3 to AF4 (blue arrow), and this is confirmed by TGCI results, which show a strong connection from C3 to AF4 (TGCI = 17.86, blue arrow). The agreement between the two methods suggests robust bidirectional dynamics during both fists’ MI, with information flow between channels over the motor and prefrontal parts.
    In both fists’ MM task, Table 7 identifies connectivity from AF3 to C3 and AF4 to C3 (red arrows). This finding is consistent with the TGCI results, which also highlight connections from AF3 to C4 (TGCI = 6.19, red arrow) and AF4 to C4 (TGCI = 5.95, red arrow). These connections emphasize the critical role of AF3 and AF4 in coordinating both fists’ movement during MM. However, the blue arrows from C3 to AF3 and C3 to AF4 in Table 7 indicate significant feedback from C3 to AF3 and AF4, which are partially captured by TGCI results, such as the connection from C4 to AF3 (TGCI = 11.9, blue arrow). This feedback highlights the bidirectional nature of the connectivity necessary for maintaining coordination during bilateral fists movement. The combination of TGC sample differences and TGCI results highlights the condition for continuous, reciprocal communication between the prefrontal and motor regions to achieve coordinated motor control.
    For both feet’ MI task, Table 7 shows significant differences from C3 to AF3 and C3 to AF4 (blue arrows). These connections indicate that C3 influences both the left and right prefrontal area channels during this task. Additionally, Table 7 reveals a significant connection between AF3 and C4 (red arrow), indicating that the AF3 influence over the C4 was evident during imagery of bilateral foot movements. Comparing these findings with the TGCI results, we observe a strong connection from C3 to AF3 (TGCI = 12.14, blue arrow) and from C3 to AF4 (TGCI = 24.05, blue arrow), reinforcing the importance of these pathways during MI. There is also a consistent outcome with regard to the connectivity between AF3 and C4 (TGCI = 6.19, as indicated by the red arrow in Table 7). The combination of both emphasizes the bidirectional dynamics necessary for coordinating complex motor imagery tasks involving both feet’.
    In the MM task for both feet’, Table 7 shows significant connectivity from AF3 to C3, AF3 to C4, and AF4 to C4 (red arrows), highlighting the prominent role of AF3 and AF4 in driving motor activity during actual bilateral foot movement. Additionally, significant feedback is observed from C3 to AF3 (blue arrow). TGCI results for both feet’ MM task further emphasize this feedback, particularly highlighting the strong connection from C3 to AF3 (TGCI = 31.67, blue arrow), which is the most pronounced feedback pathway observed. The consistency between the significant differences in TGC samples and TGCI findings suggests that both methods effectively capture the critical bidirectional flow of information between the channels over the prefrontal and motor areas during MM. The significant connections identified by TGCI (e.g., AF3 to C4 with TGCI = 19.29, red arrow and AF4 to C3 with TGCI = 11.9, red arrow) further support the argument that the AF3 and AF4 take a part in the bilateral motor activity required for both feet’ movement.
    MI vs. MM Task Comparisons: In the left-hand fist movement, Table 7 reveals an interplay between MI and MM. The green arrows from AF4 to C4 and C4 to AF4 suggest that MI engages feedback connectivity between the right prefrontal and right motor compared to MM, which is consistent with the TGCI results showing significant connectivity from AF4 to C4 (TGCI = 25.24, red arrow) and C4 to AF4 (TGCI = 13.1, red arrow) during MI. This indicates that AF4 and C4 are more strongly engaged during MI for left-hand fist movement than during MM. On the other hand, the purple arrows from AF3 to C3, from AF3 to C4, from C3 to AF3, from C3 to AF4, and from C4 to AF3 in Table 7 suggest that MM necessitates enhanced connectivity between these channels when compared to MI. This pattern aligns with the TGCI findings that emphasize the importance of feedback during actual movement, suggesting that during MM, C3, and C4 have a more prominent influence on AF3 and AF4.
    For the right-hand fist task, the significant differences in TGC samples in Table 7 show a mixed pattern of connectivity between MI and MM. The green arrows from C3 to AF4 and from C4 to AF4 suggest that MI relies more heavily on motor-to-prefrontal influence. This observation contrasts with the TGCI results, where MI shows significant prefrontal-to-motor connectivity, particularly from AF3 and AF4 to C4 (TGCI = 18.33, red arrow). Similar discordance was observed between the statistically significant differences in TGC samples and TGCI outcomes in the right-hand fist MI task. On the other hand, the purple arrows from AF3 to C4, AF4 to C3, and C4 to AF3 in Table 7 indicate that MM involves stronger prefrontal-to-motor connectivity, slightly aligns with the TGCI results that highlight AF3 to C4 (TGCI = 18.33, red arrow) and AF4 to C4 (TGCI = 6.43, red arrow) as critical connections during right-hand fist movement.
    In both fists’ movement, Table 7 column 3 shows a clear preference for MI in terms of bidirectional connectivity, as indicated by the green arrows from AF3 to C4, AF4 to C3, C3 to AF4, C4 to AF3, and C4 to AF4. These findings suggest that MI requires a more integrated and reciprocal interaction between channels over prefrontal and motor regions to coordinate bilateral fist movements. The TGCI results support this observation, with significant connectivity noted from C3 to AF4 (TGCI = 17.86, blue arrow) and from C4 to AF3 (TGCI = 11.90, blue arrow) during MI. Conversely, the absence of purple arrows in Table 7 for both fists’ movement indicates that MM does not engage these bidirectional connections as strongly as MI, suggesting a more straightforward connectivity pattern during actual movement.
    In both feet’ movement task, Table 7 highlights a strong preference for MI, as evidenced by the green arrows indicating stronger bidirectional connectivity from AF4 to C3, AF4 to C4, C3 to AF3, C3 to AF4, C4 to AF3, and C4 to AF4. These results suggest that MI of both feet’ involves complex interactions between the channels over prefrontal and motor areas, likely due to the cognitive demands of coordinating bilateral foot movements. The TGCI results corroborate this, with strong feedback from C3 to AF4 (TGCI = 24.05, blue arrow) and from C4 to AF4 (TGCI = 18.10, blue arrow) during MI. The purple arrow from AF3 to C4 in Table 7 suggests that MM involves stronger prefrontal-to-motor connectivity in this specific pathway, aligning with the TGCI result showing significant connectivity from AF3 to C4 (TGCI = 19.29, red arrow) during MM. This difference in connectivity patterns between MI and MM for both feet’ motion highlights the distinct neural circuitries employed for imagined versus actual movement, with MI requiring more extensive feedback and coordination across hemispheres.
    In summary, the comparison between MI and MM across different tasks reveals that MI generally engages in more complex and bidirectional connectivity, particularly in tasks involving bilateral coordination. This is reflected in the significant differences in TGC samples and supported by TGCI results, which together provide an important view of the distinct neural mechanisms underlying motor imagery and motor movement. The significant differences in TGC samples are particularly sensitive to bidirectional feedback dynamics, capturing refined interactions that are also reflected in TGCI results, though sometimes with different emphases. TGCI, on the other hand, provides a clear quantification of the overall strength and directionality of connectivity, highlighting key connections that drive motor control and imagery. In light of the two distinct perspectives from which these connections can be evaluated, it seems more analytical to perceive seemingly unrelated dissimilarities as complementary rather than exclusive.
    The findings of the study align and cohere with the literature (Chen et al., 2009; Vigasina et al., 2023; Drapkina et al., 2024) on MI and motor execution ME tasks, confirming established findings without fundamental contradictions. Similar studies have examined EEG-based brain connectivity during motor tasks, focusing on regions like the sensorimotor and frontal cortices. The application of TAR models here offers a novel contribution by addressing time-varying nonlinear causal relationships between EEG channels, which earlier works did not address. For instance, while previous research by Solodkin et al. (2004) has shown less complex network activation during MI compared to ME, the current study provides a more in-depth, temporally resolved view of how these connectivity patterns evolve. The focus on delay-specific interactions and the inclusion of statistical causality analysis adds new dimensions to the understanding of dynamic brain connectivity, enhancing the field’s knowledge from a fresh perspective. This study advances the current methodology by integrating temporal causality into brain network analysis, offering valuable insights into motor-related neural dynamics.
    4 Conclusion
    This study proposes the TAR4C approach as a new tool for exploring nonlinear and time-varying causal interactions in networks of time series, particularly in EEG data. The conceptual scope of TAR4C includes threshold nonlinearity detection and directional causality estimation. TAR4C, herein, offers a novel perspective on network connectivity by focusing on the detection of threshold nonlinearity and the estimation of threshold Granger Causality.
    The application of the TAR4C reveals different connectivity patterns in EEG data from MI and ME tasks. These suggest that MI engages more bidirectional and denser connectivity, especially in movements that demand coordinated action, such as opening and closing both feet. MI, which involves the mental practice of the movement without actual execution, operates more elaborate neural circuits. On the contrary, ME practices simpler connectivity patterns involving feedback mechanisms among channels.
    The findings show that bidirectional connectivity is important in MI tasks. The differences observed in TGC sampling distributions suggest that MI involves more reciprocal interactions between channels. This bidirectional flow of information appears to be important for the cognitive processes that underlie the mental rehearsal of movement, where the brain simulates both the sensory and motor aspects of an action without actually performing it.
    The TAR4C approach has its limitations. Estimating TAR models, particularly in high-dimensional networks, is a computationally burdensome process that may restrict its applicability to larger datasets or more complex networks. One way to handle larger datasets and more complex networks is to use multivariate TAR models. Multivariate TAR models extend the TAR framework to multiple time series, allowing for the simultaneous modeling of interactions among several components within a network. Furthermore, the proposed approach is applied only to EEG recordings during motor tasks, which provide valuable practical insights. Therefore, expanding TAR4C to other types of brain activity (e.g., cognition, mental disorders) or different modalities (e.g., MEG, fNIRS, fMRI) could help examine its broader applicability.
    Beyond neuroscience, TAR4C can be applied in fields dealing with complex time series data, such as economics, finance, and climate studies. For example, in economics or finance, where one needs to understand the dynamic interactions of various economic indicators or several stock prices, it can detect threshold nonlinear dependent causal connections that may remain undercover while using conventional methods. Similarly, it can be applied in climate science to investigate causal interactions at various locations between meteorological variables to let us understand how connectivity changes under varying climate conditions.
    In conclusion, TAR4C can estimate cerebral connectivity from an unexplored perspective. It can be used in many fields to understand nonlinear and time-varying interactions. The approach can become computationally demanding. However, its ability to explore complex interactions utilizing threshold nonlinearity based inference makes it valuable for neuroscience and beyond.
    SUPPLEMENTARY MATERIAL
    EEG MM/MI data set:
    The data set used in the article can be downloaded from the physionet.org repository using the link below.
    https://physionet.org/content/eegmmidb/1.0.0/