### v2

# Nonlinear Causality in Brain Networks: With Application to Motor Imagery vs Execution

Abstract
One fundamental challenge of data-driven analysis in neuroscience is modeling causal interactions and exploring the connectivity between nodes in a brain network. Various statistical methods, using different perspectives and data modalities, have been developed to understand the causal structures in brain dynamics. This study introduces a novel statistical approach, TAR4C, to dissect causal interactions in multichannel EEG recordings. TAR4C uses the threshold autoregressive (TAR) model to describe causal interactions between nodes in a brain network from two perspectives. The first tests whether one node controls the dynamics of another. The controlling node, named the threshold variable, implies its causative role since it operates as a switching mechanism governing the instantaneous transitions between autoregressive structures. This concept is known as threshold non-linearity. Once verified between a node pair, the next step in TAR modeling is assessing the causal node’s predictive ability on the other’s activity, representing causal interactions in autoregressive terms, a concept underlying Granger (G) causality. TAR4C can discover non-linear, time-dependent causal interactions while maintaining the G-causality framework. The approach’s efficacy is demonstrated through EEG data from a motor execution/imagery experiment. By comparing causal interactions during motor execution and imagery, TAR4C reveals key similarities and differences in brain connectivity across subjects.
Keywords: Regime Switching Time Series Models; Threshold AR; Nonlinear Causality; Time-varying Causality; Brain Connectivity
1 Introduction
The exploration of the complex connectivity and causality intrinsic to networks of time series data represents a pivotal challenge that necessitates significant methodological enhancements, particularly within the domain of neuroscience (Biswas and Shlizerman, 2022). Many experiments over the last few decades have produced increasing evidence that brain connectivity is both dynamic and nonlinear, thus demanding sophisticated analytical approaches to decode the underlying interactions. This study is motivated by the challenging problem of differentiating brain networks between motor imagery vs. motor execution tasks. This has a potential impact on identifying the role of brain function (in particular, connectivity dynamics) in the treatment of stroke. Our aim is to propose a conceptually comprehensive methodology that captures non-linear causal connectivity between any given nodes of interest (NOIs) within a network of time series. To align with this aim, a key analytical objective is set to explore (Granger) causal interactions within a network of electroencephalogram (EEG) time series data. While conventional and state-of-the-art methods are undoubtedly practical and helpful, they are limited in their ability to incorporate the nonlinear, especially dynamic phase-shifting (i.e., phase transition) and time-varying nature of these interactions (Sporns, 2016).
1.1 Literature Background Specific to Causal Inference in EEG
A significant number of methods have been proposed as means of inferring causality and connectivity in EEG, with varying degrees of focus on specific objective aspects, including Granger Causality (GC)(Seth et al., 2015; Ombao and Pinto, 2022), Transfer Entropy (TE) (Vicente et al., 2011; Redondo et al., 2023), Directed Transfer Function (DTF) (Babiloni et al., 2005), Partial Directed Coherence (PDC) (He et al., 2014), Bayesian modeling (Prado et al., 2001; Gorrostieta et al., 2012; Yu et al., 2016; Chiang et al., 2017), and Machine / Deep Learning (ML / DL) modeling (Shojaie and Fox, 2022)
The conventional Granger Causality method employs linear autoregressive models, which assess whether the past values of one time series can improve the prediction of another (as determined by lessening in squared prediction error). Due to the inherent linearity assumption of conventional GC, its usage in the complex, nonlinear dynamics of brain signals will lead to limited applications and interpretations (Shojaie and Fox, 2022). To overcome these restrictions, alternative approaches have been developed to permit the Granger causality framework to be extended to the nonlinear dynamics characterizing real-world contexts (e.g., Tank et al., 2022). The challenges to developing methodology for nonlinear GC include the difficulty of determining the degree of nonlinearity, the prevalence of overfitting issues, the incompatibility of the method with high dimensionality, the generation of high variance estimators, the necessity for a large number of hyperparameters, the use of black-box routines, and the complexity of interpreting the results.
On the other hand, TE is a nonparametric, model-free method for quantifying directional dependencies between time series, particularly useful for capturing nonlinear interactions. However, TE faces challenges related to computational complexity, data requirements, and interpretability. It requires large datasets for reliable estimates, and its sensitivity to noise and data sparsity, especially in EEG, can lead to false positives or spurious connections.
In contrast, machine learning models have gained popularity for their ability to detect complex, nonlinear relationships in EEG data, but they are often criticized for their lack of interpretability. These models can effectively classify and predict brain states but they often operate as ”black boxes,” offering little insight into the underlying causal mechanisms of neural activity. Moreover, machine-learning approaches require large, well-labeled datasets, which are often scarce in neuroscience. Reproducibility across studies can also be a concern, as model assumptions, dataset variability, and preprocessing differences can result in inconsistent findings across studies (Wein et al., 2021). Despite these challenges, ML’s flexibility and capacity for handling vast amounts of data offer exciting prospects, especially when combined with rigorous causal inference methods.
1.2 Challenges of Inferring Causal Interactions in EEG Data
Inferring causal interactions in brain activity, as it responds to stimuli, poses several challenges, largely due to the brain’s complex, nonlinear, and time-varying dynamics. Brain functioning is characterized by various types of nonlinearities (Breakspear, 2017) Neural nonlinearities are complex relationships in brain function that appear in various forms (e.g., synaptic plasticity, neural oscillations, neuronal firing rates). This paper aims to analyze the abrupt phase-shifting dynamic of the human brain from a previously unexplored and undiscussed nonlinearity perspective in relation to EEG data. The approach we propose is in line with the findings that the brain operates with nonlinearity characterized by abrupt phase shifts when processing information.
A further challenge in inferring causality from EEG data resides in the limited capacity of EEG to depict spatial resolution of neuronal activities in sufficient detail. Thus, neural sources are not localized complicating the task of identifying genuine causal relationships between brain regions. The recorded signals often reflect aggregated activity from multiple sources, distorted by volume conduction, which can create undecomposable correlations and spurious causal inferences.
Source localization and beamforming can help mitigate these issues but they also introduce their own limitations and assumptions that must be carefully considered. Moreover, the causal inference derived from EEG often reflects a simplified model of brain connectivity. While EEG reveals the directional flow of information between channels, it offers limited insight into the biophysical processes involved, such as synaptic or neurotransmitter activity. Therefore, the causal relationships inferred from EEG data should be interpreted with caution, as they represent statistical associations rather than direct neural interactions (Bastos and Schoffelen, 2016; Ombao and Pinto, 2022).
1.3 Objective of this Paper
We develop a new approach, the Threshold Autoregressive model for Causality (TAR4C), to discover nonlinear and time-varying causal interactions in a high dimensional network. The TAR4C method minds the gap in existing methods, which can only comprehend several manifestations of nonlinearities. By using a threshold nonlinearity perspective, TAR4C offers a way to explore brain dynamics, particularly phase transitions in brain connectivity, which reflect how a reference channel may impact the interaction between nodes. The specific objectives are as follows:
Introducing Threshold Nonlinearity for Connectivity Inference. The existing methods fail to consider the operating principle of the brain, which is characterized by abrupt transition dynamics, and tend to prioritize particular facets of nonlinearity. The threshold variable (i.e., node/channel) in the TAR framework serves as a triggering reference for the other node. In other words, the level of activity in the threshold variable governs the other variable’s activity. It can thus be reasoned that the two nodes (e.g., regions) are connected in this respect alone. This approach identifies phase transitions based on threshold (e.g., cause) variables, providing a new way to understand the brain’s connectivity and dynamic processes.
Modeling Nonlinear and Time-Varying Causal Interactions. Another goal is to model how different nodes in a network influence each other under different conditions by validating nonlinear and time-varying interactions. Applying TAR4C to EEG data enables us to study and test the causal relationships as threshold variable’s states change, particularly during motor tasks like execution and imagery. TAR4C is, therefore, capable of tracking causal connectivity alterations over time, which is not addressed by current methodologies.
Comparing Connectivity Patterns in Motor Execution and Imagery. The study also aims to examine differences in brain connectivity during motor execution (ME) and motor imagery (MI). It investigates how causal interactions between regions, such as the prefrontal and motor cortices, vary across tasks involving different gestures (e.g., open and close of both fists, both feet). By applying TAR4C, the study discovers distinct patterns in how the brain responds to real and imagined motor tasks.
Quantifying Connectivity with New Indices. The TAR4C method introduces two new metrics: the Threshold Connectivity Index (TCI) and the Threshold Granger Causality Index (TGCI). These indices measure the strength and direction of causal interactions between EEG channels (e.g., nodes). The TCI identifies threshold-based connectivity, while the TGCI evaluates the predictive relationships between channels. These indices provide a simplified representation of nonlinear brain interactions.
Expanding TAR4C Beyond EEG Networks. Although the data analytic focus is on EEG data, the TAR4C framework can also be applied to other networks of time series, such as those pertaining to economics or climate variables, especially when the objective is to discover relationships between specific variables in a high-dimensional setting. This study establishes TAR4C as a tool for detecting nonlinear causal interactions across various domains, highlighting the broader relevance of threshold autoregressive models in understanding complex connectivity patterns.
2 Exploring Connectivity and Causality via Threshold Autoregressive Models
This study proposes an approach based on threshold autoregressive (TAR) modeling to explore connectivity within time series networks and a novel application to multi-channel motor EEG recordings presented as a case study. EEG devices measure brain activity through signals captured by electrodes (channels/nodes) on the scalp. These signals, rich in temporal information, are valuable data for studying brain connectivity during tasks (e.g., motor). The brain, a highly dynamic structure, constantly shifts between different connectivity states in response to stimuli and tasks. This study assumes that phase transitions are governed by phase-shifting mechanisms that form effective connections between brain regions. The TAR model, designed to handle regime shifts, can track these transitions by specifying functionals that control the dynamics. By identifying these phase changes, the TAR model provides a practical tool for discovering causal interactions among EEG channels, offering insights into how different channels influence one another.
2.1 Inference in high-dimensional networks of time series
Most phenomena in the real world are composed of many interconnected components. Statistical inference in such high-dimensional networks of time series continues to pose a significant analytical challenge. In such interconnected multi-node networks, if the primary purpose and scope of the research are to determine the connectivity between any specific two nodes, the confounding effects of the other nodes in the network cannot be ignored. Since the estimation and interpretation of multi-parameter models become more difficult as the number of parameters increases, instead of using a multi-parameter model, we adopt a practical and innovative dimensionality reduction approach proposed by Aslan and Ombao (2024) to determine the connectivity between two specific nodes accounting for the effect of all the other nodes in the network. This approach employs spectral dynamic principal component analysis (sDPCA) in the frequency domain to create a low-dimensional representation of all other nodes. This approach was originally developed in Brillinger (1974). The resulting principal component scores are subsequently removed from the nodes of interest, thus eliminating the confounding effects. This method ensures that the effects of other nodes are effectively isolated, allowing for a precise determination of connectivity between the two nodes.
Let $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ be a network of time series with $n$ nodes. In this paper, we consider a network representing brain activity through EEG recordings, where each component (i.e., node or channel) is a time series. Within this network $N$, we let $Y^{*}$, and $X^{*}$ represent any two nodes of interest. As demonstrated in Aslan and Ombao (2024), sDPCA in the frequency domain offers practical benefits and is highly effective for the alleviation of dimensionality problems in high-dimensional settings. Thereby, the interference-free counterparts of the nodes of interest can be obtained in the following manner:
$Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
$X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|f_{z}(Z_{1,t},\ldots,Z_{n-2,t})\big{)},$
where $f_{z}(.)$ is a function that accounts for the dynamic variability of the excluded nodes. In practice, $f_{z}(.)$ serves as a filter and can take any form, provided it effectively summarizes the net discharge of the excluded nodes.
2.2 Background on the threshold autoregressive model
The concept of switching between regimes based on a change point was first introduced by Quandt (1958) in linear regression. Two decades later, Tong (1978) and Tong and Lim (1980) extended this idea to nonlinear time-series analysis through threshold non-linearity. Since then, the threshold autoregressive (TAR) model has been applied widely across various domains. TAR models offer practical methodologies for exploring complex dynamical systems in time series, as they capture nonlinear dynamics, limit cycles, abrupt jumps, and asymmetries. The primary purpose of the TAR model is to identify threshold-driven regimes in time-varying systems and provide autoregressive representations of these regimes. The thresholding mechanism makes the TAR model particularly useful, as it decomposes complex stochastic systems into simpler subsystems, allowing for more efficient implementation and better interpretation. TAR models are well-suited to represent systems with shifting levels and sudden bursts, often seen in time series data such as EEG signals. For a comprehensive review of the applications of threshold nonlinearity and TAR models, we refer the reader to Hansen (2011) and Chen et al. (2011). We introduce the TAR model as an effective analytical framework in this study to gain insight into the intricate and highly dynamic brain activity via EEG signals.
A stochastic process $Y_{t}$ is TAR with $k$ regimes if it can be represented as follows
$Y_{t}=\sum_{j=1}^{k}\left[(\phi_{0}^{(j)}+\sum_{i=1}^{p_{j}}\phi_{i}^{(j)}Y_{t-i})\ I(r_{j-1}<Z_{t-d}\leq r_{j})\right]+\ \varepsilon_{t},$ (1)
[…]
$Y_{t}=\phi_{0}^{(j)}+\sum_{i=1}^{p_{j}}\phi_{i}^{(j)}Y_{t-i}\ +\ \varepsilon_{t}^{(j)}.$
The overall process $Y_{t}$ is nonlinear if there are at least two regimes with different models, even if each is linear.
The unknown parameters for the model in Equation (1) can be assembled as follows
$\Omega=\left[(\phi_{0}^{(1)},\phi_{i:p_{1}}^{(1)}),\ldots,(\phi_{0}^{(k)},\phi_{i:p_{k}}^{(k)})\right].$
The hyperparameters of the model can be shown as;
$\Gamma=\{k,p_{j},d,r\in[\text{min}(Z_{t-d}),\text{max}(Z_{t-d})]\}.$
Since the TAR model is a piece-wise linear model, the estimates of the parameters can be readily obtained by using the least squares (LS) method conditional on hyperparameters (i.e., estimation is equivalent to the maximum likelihood estimation under the auxiliary assumption that $\varepsilon_{t}$ is i.i.d.$N(0,\sigma_{\varepsilon_{t}}^{2})$). Thus, the conditional least squares (CLS) method has a significant advantage in that it allows the exploration of the threshold parameter and hyperparameters without imposing any restrictions. The consistency of the CLS method was demonstrated in Chan (1993) to hold under certain mild conditions such as stationarity and the presence of a discontinuous conditional mean function.
The LS estimators $\widehat{\Omega}$, $\widehat{\Gamma}$ jointly minimize the following objective function:
$\mathcal{S}({\Omega},{\Gamma})=\sum_{t=1}^{T}\left[Y_{t}-\sum_{j=1}^{k}\left[(\phi_{0}^{(j)}+\sum_{i=1}^{p_{j}}\phi_{i}^{(j)}Y_{t-i})I(r_{j-1}<Z_{t-d}\leq r_{j})\right]\right]^{2},$ (2)
$\widehat{\Omega},\widehat{\Gamma}=\operatorname*{\arg\min}_{\Omega,\Gamma}\mathcal{S}({\Omega},{\Gamma})$ (3)
where $T$ denotes the sample size. The LS estimators $\widehat{\Omega}$ and the minimization of the objective function in Equation (2) can be achieved conditionally by a grid search considering all viable combinations of hyperparameters $\{j,p_{j},r,d|j=1,\ldots,k\}$. The optimization requires the execution of approximately $T^{k-1}/2\times d\times p_{(1)}\times\ldots\times p_{(k)}$ arranged autoregressions. Arranged autoregression is an autoregression where cases are rearranged by $\Gamma$-dependent sample splitting. Note that rearrangement of the sample based on the threshold values does not change the temporal structure of the series within cases (Tsay, 1989). The asymptotic theory of LS estimation and inference for TAR models was addressed and developed by Chan (1993) and Zhang et al. (2023) (among many others).
Remark. The model in Equation (1) is called self-exciting threshold autoregressive (SETAR) model if the threshold variable is the same as the actual time series, i.e., $Z_{t-d}=Y_{t-d}$. In this case, the regime of the time series $Y_{t}$ would be determined by its own past value, $Y_{t-d}$.
2.3 Proposed Approach: TAR for causality in EEG networks
EEG functions as a network of interconnected components, where connectivity between node pairs can be the subject of investigation. This study proposes a 2-regime TAR model based approach to dissect causal interactions between EEG channels. The main objective is to map the net causal influence each node receives from others through pairwise connectivity analysis while minimizing interference from excluded nodes. Let $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$ represents $n$-dimensional network of multiple EEG time series; $Y^{*}$ and $X^{*}$ be the nodes of interest in this network, between which we aim to discover the causal relationship. Considering $Y$ and $X$ be the neutralized counterparts of $Y^{*}$ and $X^{*}$ from the impact of other nodes in the network, the ”net” activities of nodes $Y^{*}$ and $X^{*}$ at time $t$ can be denoted as $Y_{t}$ and $X_{t}$. Thereby, the 2-regime TAR model for the node $Y$, and the node $X$, which is treated as both the threshold and predictive variable, can be formulated as follows:
$Y_{t}=\begin{cases}\phi_{0}^{(1)}+\sum_{i=1}^{p_{1}}\phi_{i}^{(1)}Y_{t-i}+\sum_{\ell=1}^{q_{1}}\theta_{\ell}^{(1)}X_{t-\ell}\ +\ \varepsilon_{t},&\text{if }u(X_{t-d})\leq r\\ \phi_{0}^{(2)}+\sum_{i=1}^{p_{2}}\phi_{i}^{(1)}Y_{t-i}+\sum_{\ell=1}^{q_{2}}\theta_{\ell}^{(2)}X_{t-\ell}\ +\ \varepsilon_{t},&\text{if }u(X_{t-d})>r\end{cases},$ (4)
where $u(X_{t-d})$ is the function constructs the threshold variable; $d$ denotes delay parameter with positive integer that establishes the time delay at which the threshold function $u(X_{t-d})$ significantly influences the structure of $Y_{t}$; $r\in[\text{min }u(X_{t-d}),\text{max }u(X_{t-d})]$ denotes threshold value; $p_{1}$ and $p_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $Y_{t}$, respectively; $q_{1}$ and $q_{2}$ indicates $1^{th}$ and $2^{nd}$ regime lag order of the node $X_{t}$, respectively; $\{\phi_{0}^{(1)};\phi_{i}^{(1)}|i=1,2,\ldots,p_{1};\theta_{\ell}^{(1)}|\ell=1,2,\ldots,q_{1}\}$ are parameters for $1^{th}$ regime; and $\{\phi_{0}^{(2)};\phi_{i}^{(2)}|i=1,2,\ldots,p_{2};\theta_{\ell}^{(2)}|\ell=1,2,\ldots,q_{2}\}$ are parameters for $2^{nd}$ regime; $\{\varepsilon_{t}\}$ is a sequence of martingale differences satisfying ${\mathbb{E}}(\varepsilon_{t}|F_{t-1})=0$, ${sup}_{t}{\mathbb{E}}(|\varepsilon_{t}|^{\delta}|F_{t-1})<\infty$ $almost$ $surely$ for some $\delta>2$, where $F_{t-1}$ is the $\sigma$-field generated by $\{\varepsilon_{t-i}|i=1,2\ldots\}$.
The unknown set of parameters for the model in Equation (4) can be shown as follows:
$\Omega=\left\{\Phi,\Theta\right\},\quad\Phi=\left\{\phi_{0}^{(1)},\phi_{i:p_{1}}^{(1)},\phi_{0}^{(2)},\phi_{i:p_{2}}^{(2)}\right\},\quad\Theta=\left\{\theta_{\ell:q_{1}}^{(1)},\theta_{\ell:q_{2}}^{(2)}\right\}.$ (5)
The hyperparameters of the model can be collected as
$\Gamma=\{p_{1},p_{2},q_{1},q_{2},d,r\in[\text{min }u(X_{t-d}),\text{max }u(X_{t-d})]\}.$
The parameters of the model in (4) can be estimated by LS method conditional on hyperparameters similar to minimization of the objective function shown in (2) and (3). To discover TAR model-based causative interactions in a network $N$ consisting of multiple sources of information, called nodes, it is helpful to specify the following definitions regarding (4) and (5).
Definition 1
[…]
(X is threshold-connected to Y if X exerts a threshold nonlinearity effect on Y)
The first definition implies that the presence of a threshold non-linearity in the TAR model (4) would be sufficient to show that $X_{t}$ has a threshold conditional causal effect on $Y_{t}$. Note that, as per this definition, the set of parameters $\Theta$ is considered an empty set. In other words, even if the past of $X_{t}$ does not have a significant predictive effect, which would be provided by $\Theta$, the threshold functional of $X_{t-d}$ has a particular influence on the structure of $Y_{t}$ in the sense that it controls which regime $Y_{t}$ will follow. To explore the threshold-connectivity (TC) stated in definition 1, we utilize the threshold non-linearity test scheme suggested by Hansen (1996, 2000).
The model given in (4) can also be rewritten to illustrate Hansen’s test in a simpler way. Let $W_{t}=(1\text{ }Y_{t-1}\text{ }\ldots\text{ }Y_{t-p}\text{ }X_{t-1}\text{ }\ldots\text{ }X_{t-q})^{\prime}$ be a $m\times 1$ vector where $p_{1}=p_{2}=p$; $q_{1}=q_{2}=q\text{ }\text{ }\text{ }\text{ }$ and $m=p+q+1$; $\pi_{1}=(\phi_{0}^{(1)}\text{ }\phi_{1}^{(1)}\text{ }\ldots\text{ }\phi_{p}^{(1)}\text{ }\theta_{1}^{(1)}\text{ }\ldots\text{ }\theta_{q}^{(1)})^{\prime}$ and $\pi_{2}=(\phi_{0}^{(2)}\text{ }\phi_{1}^{(2)}\text{ }\ldots\text{ }\phi_{p}^{(2)}\text{ }\theta_{1}^{(2)}\text{ }\ldots\text{ }\theta_{q}^{(2)})^{\prime}$ then the 2 regime TAR model in (4) takes the form;
M2: $Y_{t}=\pi_{1}^{\prime}W_{t}I(u(X_{t-d})\leq r)+\pi_{2}^{\prime}W_{t}I(u(X_{t-d})>r)+\varepsilon_{t}.\phantom{AAAAAA}$ (6)
In the absence of threshold nonlinearity, $\pi_{1}=\pi_{2}$ (i.e., $r=r_{0}$) the linear autoregressive distributed lag (ADL) model can be expressed as follows:
M1: $Y_{t}=\pi_{1}^{\prime}W_{t}+\varepsilon_{t}.\phantom{AAAAAAAAAAAAAAAAAAAAAAAAAAAA}$ (7)
The unknown parameters of the models represented in (6) and (7) can be shown as $\Pi_{\text{\tiny M2}}=[\pi_{1},\text{ }\pi_{2},\text{ }r,\text{ }d]$ and $\Pi_{\text{\tiny M1}}=[\pi_{1}]$, respectively. Under the assumption ${\mathbb{E}}(\varepsilon_{t}|F_{t-1})=0$ the appropriate estimation method is LS method for both. The LS estimator of $\widehat{\Pi}_{\text{\tiny M1}}$ and $\widehat{\Pi}_{\text{\tiny M2}}$, which is conditional on fixed hyperparameters $\{r,d\}$, can be found by the minimization of following objective functions;
$\widehat{\Pi}_{\text{\tiny M1}}=\operatorname*{\arg\min}_{\Pi_{\text{M1}}}\sum_{t=1}^{T}\big{[}Y_{t}-\pi_{1}^{\prime}W_{t}\big{]}^{2}$ (8)
$\widehat{\Pi}_{\text{\tiny M2}}=\operatorname*{\arg\min}_{\Pi_{\text{M2}}}\sum_{t=1}^{T}\big{[}Y_{t}-\pi_{1}^{\prime}W_{t}I(u(X_{t-d})\leq r)-\pi_{2}^{\prime}W_{t}I(u(X_{t-d})>r)\big{]}^{2}$ (9)
Models given in (6) and (7), which are the most restricted, have a nested structure. Hence, comparing them using a model selection criterion is equivalent to a likelihood ratio test (Hansen, 1999).
Let $S_{\text{\tiny M1}}={\hat{\varepsilon}_{\text{\tiny M1}}}^{\prime}\times{\hat{\varepsilon}_{\text{\tiny M1}}}^{\text{}}$ and $S_{\text{\tiny M2}}=\hat{\varepsilon}^{\prime}_{\text{\tiny M2}}\times{\hat{\varepsilon}_{\text{\tiny M2}}}^{\text{}}$ denote the sum of squared residuals obtained by LS estimation in (8) and (9), respectively; and let $r_{0}$ is denoting the true value of the threshold value $r$; then testing the existence of threshold non-linearity can be accomplished by the testing the hypothesis of $\text{H}_{0}:r=r_{0}$ considering the following likelihood ratio (LR) test statistic;
$LR(r)=T\bigg{(}\frac{S_{\text{\tiny M1}}-S_{\text{\tiny M2}}(\hat{r})}{S_{\text{\tiny M2}}(\hat{r})}\bigg{)}.$ (10)
The test of the hypothesis of $\text{H}_{0}$ (i.e., M1 against M2 or linear vs. non-linear model) is to reject for large values of $LR(r_{0})$. This is the LR test under the auxiliary assumption that $\varepsilon_{t}$ are independent $N(0,\sigma^{2})$. It also corresponds to the conventional F (or Wald) and the Lagrange multiplier (or score) tests. However, the asymptotic distribution of the $LR$ does not follow a chi-square distribution since the nuisance parameter $r$ is not identified and does not enter the model under the null hypothesis of no threshold effect. To overcome this barrier, Hansen (1996) demonstrates that the asymptotic distribution and thus $p$-values can be approximated by a residual bootstrap simulation. The test procedure also provides a heteroskedasticity-consistent Lagrange multiplier test that adjusts for heteroskedastic error terms.
The asymptotic distribution of the threshold nonlinearity test and some of the major theoretical problems of inference for TAR model estimation are comprehensively studied by Hansen (1996, 1997, 1999, 2000). Through testing the existence of threshold non-linearity, we aim to discern the connectivity between channels, shedding light on the directed connectivity within the brain.
Definition 2
[…]
(X threshold Granger causes Y if X has both threshold nonlinearity effect on Y and also having predictive power on Y). The second definition specifies the threshold Granger causality (TGC) by incorporating the impact of the threshold node’s predictive ability in the model. Once the threshold nonlinearity is verified, the form of the causal interactions between the two nodes can be explored and represented in terms of the TAR model. In this second definition, the threshold node $X_{t-d}$, whose threshold connectivity to $Y_{t}$ is determined by the threshold non-linearity test, is included as a predictor of the current state of $Y_{t}$. Hence, the identification of TGC would require the presence of statistically significant coefficients for any of the time-lagged terms of the $X_{t}$. Given the possibility of multiple valid representations of causal interactions, it is necessary to establish a model selection criterion that avoids over-parameterization for a well-refined representation of the causal connections. To this end, we suggest employing the heteroscedasticity-robust Wald statistics provided by Hansen (1997) to be used as a term to be penalized in model selection criteria. The Wald test for causality can be structured to test the predictive capacity of the threshold variable included as explanatory covariates in the model. It should be noted that Li (2006) was the first to mention the TGC nomenclature and made use of the Wald test, proposed and studied in detail by Hansen (1996, 1997) for the TAR model specification (i.e., testing the restricted vs unrestricted models). In his model, Li (2006) utilized the dependent variable itself as the threshold variable. Therefore, it is worth mentioning that whilst we adopt the common TGC nomenclature, our suggested methodology diverges in terms of the causal connectivity framework.
Considering the model given in (4) the hypothesis to be tested for TGC can be expressed as follows:
$\text{H}_{0}:\theta_{1}^{(1)}=\theta_{1}^{(2)}=\ldots=\theta_{q_{1}}^{(1)}=\theta_{q_{2}}^{(2)}=0,$ (11)
where the alternative hypothesis discerns the TGC by stating that at least one coefficient is nonzero. If the threshold value $r$ splitting the sample and delay $d$ are known, the pointwise heteroscedasticity-consistent Wald statistic can be shown as follows:
$\Delta_{T}(r,d)=\bigg{(}\text{C}\hat{\Omega}(r,d)\bigg{)}^{\prime}\bigg{[}\text{C}\bigg{(}M(r,d)^{-1}V(r,d)M(r,d)^{-1}\bigg{)}\text{C}^{\prime}\bigg{]}^{-1}\bigg{(}\text{C}\hat{\Omega}(r,d)\bigg{)},$ (12)
where C is the selector matrix; $M(r,d)=\sum_{t=1}^{T}W_{t}(r,d)W_{t}(r,d)^{\prime}$;
$V(r,d)=\sum_{t=1}^{T}W_{t}(r,d)W_{t}(r,d)^{\prime}\hat{\varepsilon_{t}}^{2}$; $W_{t}(r,d)=\bigg{(}W_{t}^{\prime}I(u(X_{t-d})\leq r)\text{ }\text{ }W_{t}^{\prime}I(u(X_{t-d})>r)\bigg{)}^{\prime}$, then according Hansen (1996); Davies (1987) the appropriate testing of $\text{H}_{0}$ can be done by using following LR statistic
$\Delta_{T}=\sup_{(r,d)\in\Gamma}\Delta_{T}(r,d).$ (13)
In contrast to the test statistic given in (10), this testing treats $(r,d)$ as fixed in both the null and alternative hypotheses, which makes it standard testing procedure, and the asymptotic distribution of $\Delta_{T}$ is $\chi^{2}_{\rho(C)}$, where ${\rho(C)}$ is the number of restrictions. The test rejects the null of the restricted model for the large values of $\Delta_{T}$. It should be noted that the restricted model tested via the hypothesis given in (11) is not the sole hypothesis implying TGC. All alternatives to the restricted models stated in the null hypothesis with respect to the $\Theta$ would imply TGC. In fact, the restricted models to be tested that indicate TGC should be investigated through a grid search since the statistic $\Delta_{T}$ varies over hyperparameters $p_{1},p_{2},q_{1},q_{2}$. Further, $\Delta_{T}$ provides the deviance between the restricted vs. unrestricted model, which makes it a proper tool to employ in model selection criteria for selecting a simpler one among statistically valid models. Thereby, following the confirmation of threshold nonlinearity, to ascertain a parsimonious manifestation of the causal interaction between $X$ and $Y$, we propose a model selection criterion, which is a modified version of the Schwarz criterion (SC),
${mSC}_{T}=\Delta_{T}-{2log(T)K},$ (14)
with $K$ is the number of parameters in the model. The model selection criterion based estimation of TGC can be defined as the maximization of the objective function given in (14),
$\max_{\Gamma}\text{ }{mSC}_{T}.$ (15)
The maximization of (14) yields estimates of parameters $\hat{\Omega}$ and $\hat{\Gamma}$ as a by-product. The use of a method of model selection for inference with a criterion analogous to that in (14) and related remarks have been given by Raftery (1995), Kass and Raftery (1995), Liu et al. (1997), and Gonzalo and Pitarakis (2002).
The whole procedure referred to as TAR4C can be simplified as follows:
1
2
Network : $N=\{Y^{*}_{t},X^{*}_{t},Z_{1,t},\ldots,Z_{n-2,t}\}$
Scope : $\text{Analyze}\text{ }{\text{TC}}_{X\to Y}\text{ and}\text{ }{\text{TGC}}_{X\to Y}\text{ }\text{ }$
Input : $X,Y\in N$ (nodes of interest in a network, $N$)
Parameters : $\Omega,\Gamma$
Data : $\{({X}_{t},{Y}_{t})\}_{t=1}^{T}$ the set of obs. of two nodes from the network $N$;
$u({X}_{t-d})$ threshold functional
Decision : (a) ${\text{TC}}_{X\to Y}$ (b) ${\text{TGC}}_{X\to Y}$ (c) ${\text{TC}}_{X\not\to Y}$
3
4
  1. 5Step 1.
    Neutralize interference from other nodes:
  1. Step 1.1.
    $Y_{t}=Y^{*}_{t}-E\big{(}Y^{*}_{t}|H(Z_{1,t},\ldots,Z_{N-2,t})\big{)}$
  2. Step 1.2.
    $X_{t}=X^{*}_{t}-E\big{(}X^{*}_{t}|H(Z_{1,t},\ldots,Z_{N-2,t})\big{)}$
  3. $H(.)$ denotes for spectral domain dynamic PCA.
  • 6Step 2.
    Testing threshold effect:
    1. Step 2.1.
      Specify the $u(X_{t-d})$ of the node $X$ whose causative effect on $Y$ is of interest.
    2. Step 2.2.
      Regarding the model in Equation (6), test the following hypothesis of linearity.
      $\text{H}_{0}:\pi_{1}=\pi_{2}$
    3. Step 2.3.
      Any significant evidence against the hypothesis of linearity indicates ${\text{TC}}_{X\to Y}$. If the decision is in favor of threshold non-linearity, proceed to the next step. Otherwise, or for different causality assessments, repeat Step 2 according to different threshold functionals.
    4. Step 2.4.
      If all else fails, the decision ${\text{TC}}_{X\not\to Y}$ can be imposed.
  • 7Step 3.
    Identify causal representation by TAR:
  • 8
    1. Step 3.1.
      $\max_{\Gamma}\text{ }{mSC}_{T}$
    2. Step 3.2.
      Obtain $\hat{\Omega}$ and $\hat{\Gamma}$ as a by-product of the maximization.
    3. Step 3.3.
      ${\text{TGC}}_{X\to Y}$ if $\hat{\Theta}\neq\{\emptyset\}$
    9
    Algorithm 1 TAR4C$(X,Y)$
    2.4 On the role of TAR models for exploring causal connectivity in a network of multiple nodes
    Unlike conventional Granger causality tests, which rely on linear assumptions, the use of TAR models allows for the detection of nonlinear causal relationships. This is particularly important as many real-world relationships exhibit nonlinearities, and detecting such relationships enhances the scope of causality analysis. The inclusion of a threshold nonlinearity test helps identify points where causal connections between nodes change qualitatively, enhancing interpretability in dynamic networks like EEG. Identifying regime-switching dynamics adds a layer of interpretability about the nature of causal impact in a dynamic network such as the EEG setting and can offer insights into the conditions under which, for example, the node $X$ influences the node $Y$. TAR models accommodate regime-specific dynamics, offering a refined representation of causal interactions where linear models may fail. This flexibility allows experts to better model the underlying causal mechanisms. However, TAR models can be more complex, leading to overfitting or over-parameterization, especially with limited data. Estimating TAR models and identifying thresholds can also be computationally intensive, posing challenges for large networks or datasets.
    3 Analysis of Motor Imagery vs Execution EEG Data
    3.1 Dataset and the experimental procedure
    This paper examines differences in brain network dynamics between motor imagery (MI) and execution (ME) tasks. The experiment collected EEG recordings from 109 volunteers with a 64-channel setup, available publicly at PhysioNet.org (Schalk et al., 2004). Participants were asked to either execute or imagine motor tasks based on visual cues. Subjects performed gestures such as opening and closing (OC) their both fists/feet (BH/BF). EEG data were sampled at 160 Hz, with annotations for rest and activity (real or imagined). Data preprocessing removed artifacts like eye blinks and muscle activity to maintain the integrity of temporal structures. The steps included applying a high-pass filter at 0.5Hz, removing 60Hz line noise, and performing Independent Component Analysis (ICA) to isolate and eliminate artifact sources. Eye and muscle-related artifacts were identified and excluded. Despite the preprocessing, some issues (e.g., sawtooth traces, excessive blinking) remained in some recordings. We narrowed the analysis to 35 right-handed subjects (17 females, 18 males) with clean recordings. To focus on causal connections, we selected four channels (AF3, AF4, C3, C4) positioned over the prefrontal and motor cortices (Figure 1), which are likely to be active regions during voluntary movement and imagery. The connectivities explored can be outlined in Table 1.
    [Uncaptioned image] Figure 1: Channels of interest.
    Table 1: Illustration of explored connections.
    Motor Execution Motor Imagery
    BH [Uncaptioned image] [Uncaptioned image]
    BF [Uncaptioned image] [Uncaptioned image]
    3.2 Analysis using the TAR4C model
    Each subject repeated the relevant task (i.e., task1: OC of BH/BF, task2: imagery OC of BH/BF) 3 times (i.e., runs). Within these three task-oriented runs, they performed each motor execution/imagery motion (i.e., BH, BF) approximately 23 times in total. Each epoch contains about 4 seconds (i.e., 640 observations).
    For each task, 23 test results were combined following Becker (1994) to determine the presence of threshold non-linearity in the task for each subject. The method for inferring connectivity for a channel pair, called TAR4C, is described in Chapter 2. In the application step, we focused on two index metrics to summarize connectivity results for all subjects.
    3.2.1 Threshold Connectivity Index
    The Threshold Connectivity Index (TCI) quantifies, considering all the subjects, condition-specific connectivity between two EEG channels based on the detection of threshold nonlinearity in their interaction.
    The TCI is calculated as follows:
    $\text{TCI}_{XY}=\left(\frac{1}{12\times N}\sum_{d=1}^{12}\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)\cdot n^{(d)}\right)\times 100,$
    where $n^{(d)}$ is the number of subjects out of the total $N$ who reject the linearity test at delay $d$. This indicates the number of subjects exhibiting threshold nonlinearity at the specified delay. $\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)$ is an indicator function; it equals 1 when $n^{(d)}\geq 0.7\times N$, meaning that the number of subjects rejecting linearity at delay $d$ equals or exceeds 70% of the total subjects.
    In practical terms, a higher TCI value suggests stronger threshold-based connectivity between the two channels, which is visualized in network diagrams with thicker lines representing stronger connections. The analysis procedure for calculating TCI involves the following steps:
    Table 2: Outline of the implementation for TCI calculation.
    Motor Execution / Motor Imagery
[…]
    3.2.2 Threshold Granger Causality Index
    The Threshold Granger Causality Index (TGCI) extends the TCI by incorporating an additional step that tests the predictive power of the threshold variable. After establishing the presence of threshold nonlinearity, the TGCI assesses whether the past values of the threshold variable significantly predict the future values of the dependent variable using the Wald test. This index, therefore, provides a measure of directional causal influence between EEG channels. The TGCI is calculated as follows:
    $\text{TGCI}_{XY}=\left(\frac{1}{12\times N}\sum_{d=1}^{12}\mathbb{I}\left(n^{(d)}\geq 0.7\times N\right)\cdot\text{Wald}^{(d)}\right)\times 100,$
    where $\text{Wald}^{(d)}$ is the number of subjects out of the total $N$ who pass the Wald test in favor of the predictive power of the threshold variable $X$ at delay $d$. This index quantifies the strength of the causal influence from $X$to $Y$. The TGCI reflects the strength and directionality of the causal relationship between the channels and is used to visualize the directional flow of information in network diagrams. A higher TGCI value indicates a stronger causal influence from the threshold variable to the dependent variable. The analysis procedure for calculating TGCI builds on the steps used for TCI and includes the following additional steps:
    Table 3: Outline of the implementation for TGCI calculation.
    Motor Execution / Motor Imagery
    Gesture to Perform Let $X$ and $Y$ represent two nodes in the EEG network. Compare whether $X$ is Threshold G - Cause $Y$ using the proposed approach given in Section 2.3. 1. Follow the steps as outlined in the TCI calculation procedure. 2. For each delay value $d$ where threshold nonlinearity is detected, apply the Wald test to assess the predictive power of the threshold variable $X_{t-d}$. 3. Combine the Wald test results over approximately 23 epochs for each subject. 4. Calculate the TGCI by averaging the Wald test results across all subjects and normalizing by the number of delays and subjects. 5. Visualize the TGCI results to identify the direction and strength of causal influence between EEG channels.
    3.3 Connectivity at Motor Imagery
    MI both hands’ fists gesture (MI-BHs). For the MI-BHs gesture, the TCI results (Table 4) show balanced and strong connectivity between the channels over the prefrontal and the motor cortices. The TCI value for AF3 $\to$ C3 is 7.38 (red arrow), while C3 $\to$ AF3 is 6.90 (blue arrow). Meanwhile, AF3 $\to$ C4 is 14.05 (red arrow); C4 $\to$ AF3 is 13.57 (blue arrow), indicating nearly equal connectivity in both directions. The connectivity involving AF4 is similarly strong and balanced, with TCI value from AF4 $\to$ C3 to be 14.05 (red arrow) and C3 $\to$ AF4 to be 12.62 (blue arrow); AF4 $\to$ C4 is 13.81, C4 $\to$ AF4 is 7.86 (blue arrow). The TGCI values for MI-BH reveal significant contralateral causal influence in particular channels: C3 $\to$ AF4 17.86 (blue arrow), AF3 $\to$ C4 12.14 (red arrow). The connection from C3 $\to$ AF4 and having no flow in the opposite direction of itself indicate a strong causal influence from the channel over the left motor cortex to the right prefrontal cortex.
    Both Fists (BHs) Both Feet (BF)
    TCI at MI
    [Uncaptioned image] [Uncaptioned image]
    TGCI at MI
    [Uncaptioned image] [Uncaptioned image]
    Table 4: Connectivity Nets at Motor Imagery
    MI both feet gesture (MI-BF). The MI-BF gesture (Table 4) shows the highest TCI values, particularly for AF3 $\leftrightarrow$ C4: AF3 $\to$ C4 21.19 (red arrow), C4 $\to$ AF3, 20.95 (blue arrow), indicating exceptionally strong and nearly equal connectivity in both directions. The connection AF4 $\leftrightarrow$ C3 and C4 also show high TCI values: AF4 $\to$ C3 13.10 (red arrow) and C3 $\to$ AF4 13.81 (blue arrow); AF4 $\to$ C4 14.76 (red arrow), C4 $\to$ AF4 14.52 (blue arrow), indicating strong and balanced connectivity. The TGCI results further emphasize this strong connectivity, with particularly high values for the C3 $\to$ AF4 24.05 (blue arrow) and C4 $\to$ AF4 18.10 (blue arrow). These findings indicate that during the imagery of both feet task, communication between the channels over the motor and prefrontal cortices continues, with the motor regions having a significant causal influence over the prefrontal areas. This prominent bidirectional interconnectivity appears to be an important element for the appropriate coordination of bilateral feet movement.
    Comparative analysis between gestures. The MI-BF gesture demonstrates stronger connectivity than MI-BHs, particularly between AF3 and C4. TCI values in MI-BF are higher (21.19 and 20.95 vs. 14.05 and 13.57 in BHs), reflecting more intense communication during MI-BF, with an increased role for the motor cortex. MI-BF highlights motor-to-prefrontal feedback, especially from [C3, C4] $\to$ AF4, pointing to stronger motor regulation of prefrontal activity during MI-BF. MI-BF also reveals more pronounced cross-hemispheric feedback between channels, suggesting denser bilateral engagement. Both tasks display strong bidirectional connectivity, with AF3 and AF4 affecting motor channels and vice versa. MI-BF suggests stronger motor-to-prefrontal feedback and overall higher connectivity, implying a more prominent dependence on motor cortex regulation during MI-BF.
    3.4 Connectivity at Motor Execution
    ME both hand’s fists gesture (ME-BHs). For MI-BHs gesture, the TCI results (Table 5) show relatively proportional connectivity across the prefrontal and motor cortices. The TCI values from AF4 $\to$ C3 is 14.29 (red arrow), and from AF3 $\to$ C4, it is 13.33 (red arrow), suggesting noteworthy connectivity over C3 and C4. However, the TGCI results for MI-BHs gesture seem to oppose this finding, with significant causal influence observed from C4 $\to$ AF4 12.14 (blue arrow) and from C3 $\to$ AF3 11.90 (blue arrow). The contrast in terms of TCI and TGCI during MI-BHs task suggests there is a continuous exchange of information between the channels, which might be related to coordinating bilateral motor activity. The bidirectional causal connectivity indicates that both hemispheres are actively engaged, with the prefrontal and motor cortices contributing to the planning and execution of the movement.
    Both Fists (BHs) Both Feet (BF)
    TCI at ME
    [Uncaptioned image] [Uncaptioned image]
    TGCI at ME
    [Uncaptioned image] [Uncaptioned image]
    Table 5: Connectivity Nets at Motor Execution
    Motor execution - both feet gesture (ME-BF). The ME-BF gesture shows the highest TCI values (Table 5) among all gestures, particularly for C3 $\to$ AF3 19.29 (blue arrow), and AF3 $\to$ C4 14.29 (red arrow). Findings suggest that ME-BF involves strong connectivity between the prefrontal and motor cortices, especially between the left prefrontal and both motor areas. The TGCI results support this interpretation, showing very high TGCI values, especially C3 $\to$ AF3 31.67 (blue arrow). This indicates a principal causal influence of the C3 over the AF3 during ME-BF, highlighting the significant feedback mechanisms necessary for executing and adjusting bilateral feet execution. A strong influence was observed AF3 $\to$ C4 19.29 (red arrow), reflecting the prefrontal cortex’s role in the execution of this bilateral movement.
    Comparative analysis between gestures. The ME-BF gesture shows stronger connectivity than ME-BHs, particularly in motor-to-prefrontal feedback. Higher TCI values (C3 $\to$ AF3: 19.29 vs. 5.95; C4 $\to$ AF3: 13.33 vs. 6.19) indicate greater motor cortex control over prefrontal areas during ME-BF, especially in the left hemisphere. In contrast, ME-BHs show weaker motor-to-prefrontal feedback, with more prefrontal control over the motor cortex. Both ME-BHs and ME-BF gestures involve prefrontal control over motor activity, but motor-to-prefrontal feedback is stronger in ME-BF, particularly from C3 $\to$ AF3. Cross-hemispheric connectivity is significant in both gestures, but stronger in ME-BF, particularly between C3, AF4, C4, and AF3. This reflects the greater bilateral coordination needed for ME-BF compared to ME-BHs movements. Both gestures show bidirectional connectivity between prefrontal and motor cortex, with stronger motor-to-prefrontal feedback in ME-BF. The important difference is the dominant role of the motor cortex in ME-BF, especially in motor-to-prefrontal regulation, and stronger cross-hemispheric interactions during ME-BF.
    3.5 Comparative analysis of threshold Granger causality: imagery vs. execution
    We have thus far discussed the connectivity for MI and ME using TCI and TGCI values, as visualized in Tables 4 and 5. Further, we performed a permutational Hotelling’s $T^{2}$ test (Curran and Hersh (2021)) on data from 35 subjects to examine causal connectivity differences within and between MI and ME tasks. For each pair, connectivity in one direction (e.g., $X$to $Y$) was compared to the reverse direction ($Y$to $X$) within MI and ME tasks. Significant differences are shown in Table 7, with columns 1 and 2 highlighting MI and ME differences, and column 3 showing differences between MI and ME in the same direction ($X$to $Y$). Each arrow represents a statistically significant directional difference: red for prefrontal-to-motor, blue for motor-to-prefrontal, green for MI prominence, and purple for ME prominence. The comparison procedure is outlined in Table 6:
    Table 6: Outline of the Implementation for Comparative Analysis of TGC differences
    Gesture to Perform
    For all subject; Let $X$ and $Y$ represent two nodes in the EEG network. 1. Apply the threshold nonlinearity test using the TAR(12) model for each epoch and combine test results. Then, generate a 35 (i.e., number of subjects) by 1 vector for each delay $d$. (a) Retain the F-test statistics for subjects where linearity is rejected and assign 0s for others. 2. Apply the Wald test on the data of subjects who reject linearity to assess the predictive power of the threshold variable $X_{t-d}$. (a) Filter the vector of F-test statistics based on subjects who pass the Wald test. (b) Retain only those elements corresponding to subjects who pass the Wald test, forming the TGC distribution for each delay $d$. 3. Compare the TGC distributions for each pairwise channel interaction $X$to $Y$ vs $Y$to $X$ within the same task (i.e., MI or ME) using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in the directional connectivity between $X$to $Y$ vs $Y$to $X$ within the same task (i.e., MI or ME). (b) Visualize significant differences with arrows: Red arrows for significant prefrontal-to-motor connectivity. Blue arrows for significant motor-to-prefrontal connectivity. 4. Compare the TGC distributions in the same direction (e.g., $X$to $Y$) across MI and ME tasks using the permutational Hotelling’s T² test. (a) Identify statistically significant differences in connectivity between MI and ME for each channel pair. (b) Visualize significant differences with arrows: Green arrows for stronger connectivity at MI compared to ME. Purple arrows for stronger connectivity at ME compared to MI.
    Differences at MI Differences at ME Differences MIvsME
    Both Fists (BHs)
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    Both Feet (BF)
    [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
    Table 7: Threshold Granger Causality Differences for Motor Imagery (MI), Motor Execution (ME), and MI vs ME across different motor tasks
    For the MI-BHs, Table 7 highlights significant connectivity from AF3 $\to$ C4 and AF4 $\to$ C4 (red arrows), which is consistent with the TGCI results showing similar connections (TGCI = 12.14 for AF3 $\to$ C4, red arrow). This alignment between the significant differences in TGC samples and TGCI underlines the importance of AF3 and AF4 in coordinating BHs movement during MI. Table 7 also identifies significant motor-to-prefrontal feedback for MI-BHs from C3 $\to$ AF4 (blue arrow), and this is confirmed by TGCI results, which show a strong connection from C3 $\to$ AF4 (TGCI = 17.86, blue arrow). The agreement between the two methods suggests robust bidirectional dynamics during MI-BHs, with information flow between channels over the motor and prefrontal parts.
    In the ME-BHs, Table 7 identifies connectivity from AF3 $\to$ C3 and AF4 $\to$ C3 (red arrows). This finding is consistent with the TGCI results, which also highlight connections from AF3 $\to$ C4 (TGCI = 6.19, red arrow) and AF4 $\to$ C4 (TGCI = 5.95, red arrow). These connections emphasize the critical role of AF3 and AF4 in coordinating the ME-BHs. However, the blue arrows from C3 $\to$ AF3 and C3 $\to$ AF4 in Table 7 indicate significant feedback from C3 $\to$ AF3 and $\to$ AF4, which are partially captured by TGCI results, such as the connection from C4 $\to$ AF3 (TGCI = 11.9, blue arrow). This feedback highlights the bidirectional nature of the connectivity necessary for maintaining coordination during bilateral fists task. The combination of TGC sample differences and TGCI results highlights the condition for continuous, reciprocal communication between the prefrontal and motor regions to achieve coordinated motor execution.
    For the MI-BF, Table 7 shows significant differences from C3 $\to$ AF3 and C3 $\to$ AF4 (blue arrows). These connections indicate that C3 influences both the left and right prefrontal area channels during this task. Additionally, Table 7 reveals a significant connection between AF3 and C4 (red arrow), indicating that the AF3 influence over the C4 was evident during bilateral MI-BF. Comparing these findings with the TGCI results, we observe a strong connection from C3 $\to$ AF3 (TGCI = 12.14, blue arrow) and from C3 $\to$ AF4 (TGCI = 24.05, blue arrow), reinforcing the importance of these pathways during MI. There is also a consistent outcome with regard to the connectivity between AF3 and C4 (TGCI = 6.19, as indicated by the red arrow in Table 7). The combination of both emphasizes the bidirectional dynamics necessary for coordinating complex motor imagery tasks involving both feet.
    In contrast, for the ME-BF, Table 7 shows significant connectivity from AF3 $\to$ C3, AF3 $\to$ C4, and AF4 $\to$ C4 (red arrows), highlighting the prominent role of AF3 and AF4 in driving motor activity during bilateral ME-BF. Additionally, significant feedback is observed from C3 $\to$ AF3 (blue arrow). TGCI results for the ME-BF further emphasize this feedback, particularly highlighting the strong connection from C3 $\to$ AF3 (TGCI = 31.67, blue arrow), which is the most pronounced feedback pathway observed. The consistency between the significant differences in TGC samples and TGCI findings suggests that both methods effectively capture the critical bidirectional flow of information between the channels over the prefrontal and motor areas during ME. The significant connections identified by TGCI (e.g., AF3 $\to$ C4 with TGCI = 19.29, red arrow and AF4 $\to$ C3 with TGCI = 11.9, red arrow) further support the argument that the AF3 and AF4 take a part in the bilateral ME-BF.
    Comparison: imagery vs. execution
    Table 7, 1st-row 3rd-column show a clear preference in favor of MI-BHs regarding bidirectional connectivity, as indicated by the green arrows from AF3 $\to$ C4, AF4 $\to$ C3, C3 $\to$ AF4, C4 $\to$ AF3, and C4 $\to$ AF4. These findings suggest that MI requires a more integrated and reciprocal interaction between channels over prefrontal and motor regions to coordinate bilateral fists movements. The TGCI results support this observation, with significant connectivity noted from C3 $\to$ AF4 (TGCI = 17.86, blue arrow) and from C4 $\to$ AF3 (TGCI = 11.90, blue arrow) during MI. Conversely, the absence of purple arrows in Table 7 for both fists movement indicates that ME does not engage these bidirectional connections as strongly as MI, suggesting a more straightforward connectivity pattern during actual movement.
    Table 7, 2nd-row 3rd-column highlights a strong preference for MI-BF, as evidenced by the green arrows indicating stronger bidirectional connectivity from AF4 $\to$ C3, AF4 $\to$ C4, C3 $\to$ AF3, C3 $\to$ AF4, C4 $\to$ AF3, and C4 $\to$ AF4. These results suggest that MI of BF involves complex interactions between the channels, likely due to the cognitive demands of coordinating bilateral feet movement. The TGCI results corroborate this, with strong feedback from C3 $\to$ AF4 (TGCI = 24.05, blue arrow) and from C4 $\to$ AF4 (TGCI = 18.10, blue arrow) during MI. The purple arrow from AF3 $\to$ C4 in Table 7, 2nd-row 3rd-column suggests that ME involves stronger prefrontal-to-motor connectivity in this specific pathway, aligning with the TGCI result showing significant connectivity from AF3 $\to$ C4 (TGCI = 19.29, red arrow) during ME. This difference in connectivity patterns between MI and ME for BF motion tasks highlights the distinct neural circuitries employed for imagined versus execution, with MI requiring more extensive feedback and coordination across hemispheres.
    In summary, the comparison between MI and ME across different tasks reveals that MI generally engages in more complex and bidirectional connectivity, particularly in tasks involving bilateral coordination. This is reflected in the significant differences in TGC samples and supported by TGCI results, which together provide an important view of the distinct neural mechanisms underlying MI and ME. The significant differences in TGC samples are particularly sensitive to bidirectional feedback dynamics, capturing refined interactions that are also reflected in TGCI results, though sometimes with different emphases. TGCI, on the other hand, provides a clear quantification of the overall strength and directionality of connectivity, highlighting key connections that drive motor control and imagery. In light of the two distinct perspectives from which these connections can be evaluated, it seems more analytical to perceive seemingly unrelated dissimilarities as complementary rather than exclusive.
    The findings of the study align and cohere with the literature (Chen et al., 2009; Vigasina et al., 2023; Drapkina et al., 2024) on MI and motor execution ME tasks, confirming established findings without fundamental contradictions. Similar studies have examined EEG-based brain connectivity during motor tasks, focusing on regions like the sensorimotor and frontal cortices. The application of TAR models here offers a novel contribution by addressing time-varying nonlinear causal relationships between EEG channels, which earlier works did not address. For instance, while previous research by Solodkin et al. (2004) has shown less complex network activation during MI compared to ME, the current study provides a more in-depth, temporally resolved view of how these connectivity patterns evolve. The focus on delay-specific interactions and the inclusion of statistical causality analysis adds new dimensions to the understanding of dynamic brain connectivity, enhancing the field’s knowledge from a fresh perspective. This study advances the current methodology by integrating temporal causality into brain network analysis, offering valuable insights into motor-related neural dynamics.
    4 Conclusion
    This study proposes the TAR4C approach as a new tool for exploring nonlinear and time-varying causal interactions in networks of time series, particularly in EEG data. The conceptual scope of TAR4C includes threshold nonlinearity detection and directional causality estimation. TAR4C, herein, offers a novel perspective on network connectivity by focusing on the detection of threshold nonlinearity and the estimation of threshold Granger Causality.
    The application of the TAR4C reveals different connectivity patterns in EEG data from MI and ME tasks. These suggest that MI engages more bidirectional and denser connectivity, especially in movements that demand coordinated action, such as opening and closing both feet. MI, which involves the mental practice of the movement without actual execution, operates more elaborate neural circuits. On the contrary, ME exhibit simpler connectivity patterns involving feedback mechanisms among channels.
    The findings show that bidirectional connectivity is important in MI tasks. The differences observed in TGC sampling distributions suggest that MI involves more reciprocal interactions between channels. This bidirectional flow of information appears to be important for the cognitive processes that underlie the mental rehearsal of movement, where the brain simulates both the sensory and motor aspects of an action without actually performing it.
    The TAR4C approach has its limitations. Estimating TAR models, particularly in high-dimensional networks, is a computationally burdensome process that may restrict its applicability to larger datasets or more complex networks. One way to handle larger datasets and more complex networks is to use multivariate TAR models. Multivariate TAR models extend the TAR framework to multiple time series, allowing for the simultaneous modeling of interactions among several components within a network. Furthermore, the proposed approach is applied only to EEG recordings during motor tasks, which provide valuable practical insights. Therefore, expanding TAR4C to other types of brain activity (e.g., cognition, mental disorders) or different modalities (e.g., MEG, fNIRS, fMRI) could help examine its broader applicability.
    Beyond neuroscience, TAR4C can be applied in fields dealing with complex time series data, such as economics, finance, and climate studies. For example, in economics or finance, where one needs to understand the dynamic interactions of various economic indicators or several stock prices, it can detect threshold nonlinear dependent causal connections that may remain undercover while using conventional methods. Similarly, it can be applied in climate science to investigate causal interactions at various locations between meteorological variables to let us understand how connectivity changes under varying climate conditions.
    In conclusion, TAR4C can estimate cerebral connectivity from an unexplored perspective. It can be used in many fields to understand nonlinear and time-varying interactions. The approach can become computationally demanding. However, its ability to explore complex interactions utilizing threshold nonlinearity based inference makes it valuable for neuroscience and beyond.
    SUPPLEMENTARY MATERIAL
    EEG MM/MI data set:
    The data set used in the article can be downloaded from the physionet.org repository using the link below.
    https://physionet.org/content/eegmmidb/1.0.0/